### v1

# OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment

Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents OmniDRCA, a parallel speech-text foundation model based on joint autoregressive modeling, featuring dual-resolution speech representations and contrastive cross-modal alignment. Our approach processes speech and text representations in parallel while enhancing audio comprehension through contrastive alignment. Experimental results on Spoken Question Answering benchmarks demonstrate that OmniDRCA establishes new state-of-the-art (SOTA) performance among parallel joint speech-text modeling based foundation models, and achieves competitive performance compared to interleaved models. Additionally, we explore the potential of extending the framework to full-duplex conversational scenarios. [^1] Our implementation will be released at https://github.com/lxchtan/OmniDRCA.
1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2023; Yang et al., 2024) have demonstrated remarkable performance across various NLP tasks, sparking growing interest in extending their capabilities to multimodal domains encompassing vision and audio. Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges. The release of GPT-4o (OpenAI, 2024b) is a significant milestone in this field, showing the potential of LLM-based spoken dialogue systems for seamless and natural interactions with users.
LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs. Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules. Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., tone, emotion), and significant latency. Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024). E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs. However, training LLMs to generate highly intelligent speech outputs remains challenging. E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
Recent research efforts on E2E models to address the above challenges have focused on two primary directions: Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs ingest speech representations as input, produce textual responses, and simultaneously utilize the LLM’s latent representations to synthesize audio. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and autoregressively generating both text tokens and speech tokens. The key distinction lies in whether the generated speech tokens are fed back autoregressively into the LLM during the generation process. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM, ensuring temporal alignment. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. Text-Driven Speech Models prevent LLMs from perceiving generated speech tokens during inference, thereby limiting dynamic adjustment of text generation. Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials; hence, in this work, we focus on enhancing joint speech-text modeling.
In this work, we propose OmniDRCA, a novel parallel speech-text foundation model. Our architecture preserves end-to-end advantages while significantly improving speech-text semantic alignment through two key innovations: 1) Dual-Resolution Speech Representations (DRSR): DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For the LLM’s speech comprehension, we introduce a grouping mechanism, mapping 25Hz discrete audio tokens to 5Hz speech representations that align much better with textual granularity, thereby leveraging LLM’s capabilities more effectively and improving speech understanding. Combined speech-text embeddings form the assistant’s autoregressive input. The multimodal LLM (MLLM) predicts text tokens. For speech generation, latent representations are ungrouped and autoregressively processed by the speech refinement model (SRM) for individual speech token prediction. During inference, due to input-output resolution discrepancies, each MLLM inference triggers $k$ parallel SRM executions, where $k$ is the grouping factor. 2) Contrastive Cross-modal Alignment (CCA): Similarity between grouped speech representations and text representations is computed to optimize their mutual distances. While grouping aligns speech-text granularity, CCA further enhances semantic consistency between modalities before they enter the LLM and further enhances speech comprehension. Considering inference efficiency for varying output requirements in practical applications, we also design system prompts to control the output modes, including text-only output, parallel speech-text output, and parallel speech-text output after chain-of-modality (CoM) (Zhang et al., 2023a). Experimental results demonstrate the efficacy of CoM in improving performance. For reproducibility, we will release all source code and model checkpoints.
Our contributions are three-fold: 1) We propose OmniDRCA, a novel parallel speech-text foundation model that enhances speech-to-speech generation based on discrete speech tokens and LLMs, through dual-resolution speech representations and contrastive cross-modal alignment. 2) Comprehensive experiments demonstrate OmniDRCA’s superiority over the state-of-the-art (SOTA) parallel model LUCY (Gao et al., 2025) and competitive performance against SOTA interleaved models across multiple spoken question answering benchmarks. 3) We further explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, OmniDRCA-Duplex supports real-time responses to user interruptions with minimal performance degradation.
2 Related Work
Speech Tokenization. Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to its competitive semantic capabilities and support for streamlined streaming speech generation.
E2E Speech Foundation Models. Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. The contemporaneous Text-Driven model Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024) conduct parallel decoding. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. Our approach differs from Moshi with two key distinctions: (1) the $k$ tokens correspond to grouped semantic speech tokens rather than disparate acoustic tokens, and (2) our speech refinement module explicitly ingests full-sequence contextual information to enhance semantic coherence. While SLAM-Omni (Chen et al., 2024) proposes grouped semantic token prediction for speech generation, our approach designs an ungrouping mechanism and incorporates a Speech Refined Model for autoregressive generation of individual speech tokens. Another contemporaneous parallel model Kimi-Audio (KimiTeam et al., 2025) is closest to our approach. To mitigate limitations of discrete tokens in audio understanding, they introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper features, preserving both semantic and acoustic information. In contrast, our approach enhances discrete token comprehension through contrastive loss and grouping strategies, thereby improving LLM’s understanding of speech tokens while maintaining representational efficiency.
Contrastive Cross-Modal Alignment. CLIP (Radford et al., 2021) aligns image and text representations through contrastive pretraining. VILA-U (Wu et al., 2024) integrates contrastive loss and reconstruction loss during vision token pretraining to enhance feature alignment. In speech research, Ye et al. (2022) leverages multi-class N-pair contrastive loss (Sohn, 2016) for speech translation. Zheng et al. (2024) leverages InfoNCE loss (He et al., 2020) for spoken language understanding. We incorporate contrastive loss to strengthen speech-text semantic alignment. Crucially, we adopt gradient stop (Chen and He, 2021) on text embeddings during contrastive loss computation – this modification preserves speech representation learning while preventing text feature distortion.
3 Methodology
Figure 1 illustrates the architecture of OmniDRCA. The system consists of two main components: (1) a Multimodal Large Language Model (MLLM) designed to process interleaved speech-text inputs and achieve cross-modal alignment, and (2) a Speech Refined Model (SRM) specialized in high-fidelity speech synthesis. The system operates through multimodal input encoding, joint representation learning with contrastive alignment, and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce multimodal context embeddings, and decoded into synchronized speech-text responses through SRM and the language model head.
3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the speech tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency.
Figure 1: Overview of OmniDRCA. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text token prediction. Hidden representations are ungrouped and processed by the SRM through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor. The generated speech tokens are then converted to speech waveform by the speech detokenizer. We omit speech tokenizer and detokenizer for simplicity.
3.2 Multimodal Large Language Model (MLLM)
Built upon text-LLMs, the MLLM learns joint speech-text modeling to process speech or text inputs while producing parallel speech and text outputs.
Parallel Joint Speech-Text Model. Leveraging the autoregressive generation capability of MLLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the MLLM at each timestep. Their embeddings are added as model input, forming a parallel speech-text architecture. Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$. This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework. To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological components are introduced for intermodal coordination.
Inner Text Streaming. Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping. A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
$\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
[…]
Cross-modal Contrastive Alignment.
To enable cross-modal alignment, a contrastive learning objective aligns grouped speech representations with their textual counterparts by maximizing their cosine similarity and computing a contrastive loss. This joint optimization ensures coordinated positioning of both modalities within the shared representation space.
$\mathcal{L}_{\text{cont}}=-\cos(E_{\text{text}}(\mathbf{t}),E_{\text{speech}}(\mathbf{g})),$ (4)
where $E_{\text{text}}$ and $E_{\text{speech}}$ are the mean of embeddings of text tokens and grouped speech tokens between sample pairs, respectively. Different from early applications of contrastive alignment between speech and text (Zheng et al., 2024), our approach eliminates negative sample computation and prevents mode collapse through gradient stop on the text branch.
3.3 Speech Refined Model (SRM)
SRM is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the MLLM as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the MLLM’s final hidden state is mapped to group-sized embeddings via linear projection
$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[MLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (5)
and time splitting
$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}]\quad\text{where}\quad\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k},$ (6)
the resulting $\mathbf{H}$ serves as conditional input for SRM that autoregressively generates speech tokens. Following our practice on MLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. SRM is trained to maximize the conditional probability:
$\mathcal{L}_{\text{SRM}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}),$ (7)
where $s_{i}$ represents the $i$-th speech token. By optimizing this objective, SRM learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from MLLM. This design enables SRM to effectively leverage the semantic and acoustic information encoded in the hidden representations of MLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
The end-to-end training objective integrates the three losses through multi-task learning paradigm:
$\mathcal{L}_{\text{total}}=\lambda\mathcal{L}_{\text{LM}}+\mu\mathcal{L}_{\text{cont}}+\gamma\mathcal{L}_{\text{SRM}},$ (8)
$\mathcal{L}_{\text{LM}}=-\sum_{i=1}^{T}\log P(t_{i}|t_{<i},\mathbf{g}),$ (9)
where $\mathcal{L}_{\text{LM}}$ is the autoregressive loss over MLLM, $\lambda$, $\mu$, and $\gamma$ are hyperparameters.
3.4 Multimodal Interaction Patterns
Table 1: Multimodal Interaction Patterns.
Pattern Name Abbr. Modality Flow
Speech-to-Multimodal S2M Speech $\rightarrow$ Joint speech-text response
Speech-to-Text S2T Speech $\rightarrow$ Text-only response
Text-to-Multimodal T2M Text $\rightarrow$ Joint speech-text response
Text-to-Text T2T Text $\rightarrow$ Text-only response
Speech-Text Chain STC Speech $\rightarrow$ Text transcription $\rightarrow$ Text response $\rightarrow$ Multimodal response
Speech-Assisted Chain SAC Speech $\rightarrow$ Text response (agent perspective) $\rightarrow$ Multimodal response
Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
Practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Furthermore, Zhang et al. (2023a) found that the chain-of-modality (CoM) methodology can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix. For training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
4 Experiments
Table 2: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks. Baseline results are from Yao et al. (2024) except LUCY (S2) and GLM-4-Voice-Small from Gao et al. (2025) and Zeng et al. (2025) respectively. Metrics are benchmark-specific (higher is better). The Avg. score averages four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg.
GPT-4o-Realtime Closed-source 74 51.6 71.7 69.7 66.7
Step-Audio-Chat Text-Driven 130B 43 56.1 56.7 36.7 48.1
MiniCPM-o 2.6 Text-Driven 8B 51 40.0 61.0 40.2 48.1
Llama-Omni Text-Driven 8B 39 22.9 45.3 10.7 29.5
VITA-1.5 Text-Driven 7B 20 28.1 46.7 23.3 29.5
GLM-4-Voice Interleaved 9B 51 32.0 50.0 36.4 42.3
GLM-4-Voice-Small Interleaved 1.5B 5.4 18.3 4.6 9.4
LUCY (S2) Parallel 7B 25.6 54.0 22.8 34.1
Moshi Parallel 7B 24 23.8 43.7 16.7 27.1
Mini-Omni Parallel 0.5B 25 12.8 22.0 6.9 16.7
OmniDRCA Parallel 8B 45 30.0 65.0 32.9 43.2
OmniDRCA-Small Parallel 2B 38 22.1 55.3 17.9 33.3
Table 3: Speech-to-Text ($S\rightarrow T$) Generation Performance on OpenAudioBench using benchmark-specific metrics (higher is better). Baseline results are from KimiTeam et al. (2025), except LUCY (S2) and Moshi from Gao et al. (2025) and MiniCPM-o 2.6 from our experiments. Reasoning QA is removed since our model is only trained on English data. Avg. is averaged over the four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Qwen2-Audio-Chat Cascaded 7B 57.2 45.2 69.7 40.3 53.1
MiniCPM-o 2.6 Text-Driven 8B 64.1 69.2 78.0 63.0 68.5
Step-Audio-chat Text-Driven 130B 56.5 73.0 72.3 56.8 64.7
Baichuan-Omni-1.5 Interleaved 7B 59.6 58.7 74.3 55.4 62.0
GLM-4-Voice Interleaved 9B 57.9 55.4 76.0 51.8 60.3
Moshi Parallel 7B 26.6 62.3 22.8 37.2
LUCY (S2) Parallel 7B 29.3 57.3 23.3 36.6
OmniDRCA Parallel 8B 58.0 51.7 79.7 47.7 59.7
Table 4: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following (Zeng et al., 2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from (Zeng et al., 2025), except MiniCPM-o 2.6 results are from our experiments.
[…]
MiniCPM-o 2.6 Text-Driven 8B 4.18 13.31
Llama-Omni Text-Driven 8B 3.92 9.18
GLM-4-Voice Interleaved 9B 4.33 7.83
Moshi Parallel 7B 3.90 7.95
Mini-Omni Parallel 0.5B 3.17 25.28
OmniDRCA Parallel 8B 3.86 6.75
OmniDRCA-Small Parallel 2B 3.87 8.86
Table 5: Ablation study of Dual-Resolution Grouping Strategy Conducted on OmniDCRA-Small with Llama Questions Dataset.
Modal Group Size S2T T2M (T/S) T2T
S2T+T2T (Understanding) 1 55.33 61.33
5 62.33 69.67
T2M+T2T (Generation) 1 51.67 / 41.00 58.33
5 42.00 / 25.33 64.00
Table 6: Ablation study of Contrastive Cross-modal Alignment (Cont), SRM, and Multimodal Architecture Design (MD) on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
Model S2M (T/S) S2T T2M (T/S) T2T STC (T/S) SAC (T/S) SUC (T/S)
OmniDRCA-Small 64.33 / 55.33 64.67 69.67 / 58.67 69.67 67.33 / 59.67 65.67 / 57.00 67.67 / 57.00
w/o Cont 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
   w/o SRM 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
   w/o MD 58.00 / 49.00 58.00 69.33 / 55.00 68.33
4.1 Experimental Setup
Datasets. Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for SRM pre-training. 17.8K hours of English speech conversations data synthesized from text are adopted for OmniDRCA training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^2] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following prior works (Yao et al., 2024; Li et al., 2025), we evaluate the performance on the widely used Spoken Question Answering benchmarks, including AlpacaEval (Li et al., 2023), Web Questions (Berant et al., 2013), Llama Questions (Nachmani et al., 2024), and TriviaQA (Joshi et al., 2017). To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, we adopt datasets from two established evaluation frameworks: UltraEval-Audio  [^3] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench  [^4] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. As discussed in Huang et al. (2025); Chen et al. (2025), the results on TriviaQA are for reference only and marked with “*” due to the absence of an official test set. Moreover, we detect performance anomalies of OmniDRCA on the UltraEval-Audio TriviaQA dataset; therefore, we replace UltraEval-Audio TriviaQA dataset with OpenAudioBench TriviaQA dataset for a more meaningful evaluation  [^5] For reference, this setup yields a score of 42.6 for MiniCPM-o 2.6 while its score on the original UltraEval-Audio TriviaQA is 40.2..
Evaluation Metrics. Following the settings of UltraEval-Audio and OpenAudioBench, we utilize G-Eval (Liu et al., 2023) to evaluate model performance on AlpacaEval, and employ accuracy to assess model performance on other datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines. Representive baselines are selected to cover diverse modeling paradigms: the closed-source model GPT-4o-Realtime (OpenAI, 2024b), cascaded system Qwen2-Audio-Chat (Chu et al., 2024), Text-Driven models including Step-Audio-Chat (Huang et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni (Fang et al., 2024), and VITA-1.5 (Fu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025) and GLM-4-Voice (Zeng et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Mini-Omni (Xie and Wu, 2024a). This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details. SRM uses Qwen2.5-0.5B as its base LLM, trained following $T2M$ paradigm. OmniDRCA uses Qwen2.5-7B-Instruct as its base LLM, while OmniDRCA-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For OmniDRCA, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for OmniDRCA-Small and SRM it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A100 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For OmniDRCA, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRM pretraining, and about $20$ hours on OmniDRCA post training.
4.2 Results and Discussions
Tables 2 and 3 compare OmniDRCA with the representative and competitive baselines across all major model categories, on speech-to-speech and speech-to-text generation performance. Table 2 shows that in prior works, best Text-Driven models substantially outperform top-tier Joint Speech-Text models, because text-driven systems do not allow speech tokens to interfere with the generation process, thereby minimally impacting the LLM’s generation paradigm. Moreover, prior interleaved models often outperform parallel models, though they exhibit higher latency (e.g., 13 text tokens for first speech in GLM-4-Voice) and lack temporal alignment, complicating multistream duplex modeling. As shown in Tables 2 and 3, OmniDRCA achieves SOTA performance under the parallel paradigm, achieving a substantial 24.9% relative gain over LUCY (S2) on the average score across three Speech-to-Speech QA benchmarks and 60.4% gain over Moshi across three Speech-to-Text QA benchmarks, excluding AlpacaEval in both cases. OmniDRCA also shows competitive performance against the interleaved GLM-4-Voice and Baichuan-Omni-1.5. However, a relatively large gap still remains when compared to text-driven systems. In future work, we plan to train the model on more extensive datasets to improve the performance.
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (WER(%)), as shown in Table 4. Among models with comparable sizes, text-driven models tend to exhibit higher WER, while joint speech-text models achieve relatively lower WER. The model size also significantly impacts WER. OmniDRCA achieves the lowest WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between OmniDRCA and GLM-4-Voice suggests room for improvement in modeling prosody and preserving acoustic details.
We conduct extensive ablation study to analyze effects of critical algorithmic designs.
Dual-Resolution Grouping Strategy. Table 5 reveals the dual effects of our grouping strategy: while improving speech understanding (12.6% relative gain in S2T with group size 5) and text capabilities (13.6% relative gain in T2T), it degrades speech generation quality by 38.2% relative in T2M. On the other hand, as shown in Table 6, SRM demonstrates remarkable effectiveness in speech token generation, achieving 245.9% and 71.4% relative improvements in S2M and T2M tasks respectively. Notably, SRM also enhances speech comprehension metrics (S2T and STC-T) by 10-16% relative, suggesting that improved speech token reconstruction capability improves latent speech representations. The text-only performance (T2T) remains stable, confirming SRM’s modality-specific effectiveness without interfering with text processing pathways. Our dual-resolution architecture resolves this conflict by applying grouping only in the comprehension phase (5Hz) while using raw frame resolution (25Hz) with SRM for generation.
Contrastive Cross-modal Alignment. As shown in Table 6, the introduction of speech-text contrastive learning for sentence-level semantic alignment substantially enhances the model’s semantic understanding capabilities. While the final evaluation metric (S2S performance) shows improvement over the counterpart without contrastive loss, the enhancement exhibits modality-specific efficacy. For instance, S2T performance improves by 2.34% absolute (64.67 over 62.33), while T2T performance degrades by 4.33%, suggesting that while improving speech comprehension, the contrastive objective partially compromises the LLM’s native text understanding capacity. Our multimodal evaluation framework effectively reveals these nuanced impacts, guiding future work on mitigating text performance degradation through more balanced loss design.
Multimodal Architecture Design (MD). Our multimodal design achieves two key goals: (1) context-aware modality selection and (2) implicit data augmentation. As shown in Table 6, on the speech generation task, STC/SAC/SUC demonstrate substantial improvements over S2M. STC achieves 61.00, substantially surpassing the direct generation baseline of 55.33. This reveals the critical role of CoM while also indicating potential optimization opportunities for our S2M model. Without incorporating the three CoM variants (STC/SAC/SUC), the model fails to adopt this generation paradigm guided by system prompts. Final S2M performance shows that training without MD results in a 7.5% performance degradation, confirming the data augmentation effect of our approach.
Inspired by Wang et al. (2025), to enable full-duplex interaction, we propose a time-division multiplexing (TDM) input stream that allows OmniDRCA to accept user speech inputs during the assistant’s speech generation phase, thereby efficiently utilizing otherwise idle time slots. The resulting OmniDRCA-Duplex achieves full-duplex dialogue capabilities. The methodology and empirical validations of OmniDRCA-Duplex are detailed in Appendix.
5 Conclusions, Limitations, and Future Work
We introduce OmniDRCA, a novel parallel speech-text foundation model that integrates joint autoregressive modeling with dual-resolution speech representations and contrastive cross-modal alignment. Experiments show that OmniDRCA sets new SOTA speech-to-speech/text generation performance among parallel models and reaches competitive performance to interleaved models.
Limitations: (1) OmniDRCA’s speech generation quality remains constrained by suboptimal speech tokenization. We will explore other tokenization methods, such as CosyVoice’s speech tokenizer, to address this limitation. (2) We will extend exploration from current speech understanding and generation tasks to broader audio comprehension/synthesis tasks, and integrate visual modality.
Appendix A OmniDRCA-Duplex
As mentioned in Introduction, we explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, we introduce a novel spoken dialogue model, OmniDRCA-Duplex, which supports real-time responses to user interruptions with minimal performance degradation. In this section, we first introduce OmniDRCA-Duplex in Appendix A.1, then present the experimental results and discussions of its full-duplex conversation capabilities in Appendix A.2.
A.1 Methodology
As shown in Figure 2, compared with OmniDRCA, the overall architecture of OmniDRCA-Duplex remains largely unchanged, which is also composed of the multimodal LLM (MLLM) and the speech refinement model (SRM). The primary differences between OmniDRCA-Duplex and the original OmniDRCA model lie in two aspects. The first aspect is the organization of the input structure. OmniDRCA-Duplex organizes user speech and assistant speech alternately along the temporal axis as input, referred to as the Time-Division Multiplexing Input Stream, while OmniDRCA adopts turn-based speech as the same as text-based LLM. For the second aspect, to enhance the model’s sensitivity to user interruptions, an additional assistant state prediction head is introduced in OmniDRCA-Duplex to determine whether muting should be activated. The following paragraphs provide detailed descriptions of these two designs.
Figure 2: The architecture of OmniDRCA-Duplex.
Time-Division Multiplexing Input Stream.
Full-duplex spoken dialogue models allow simultaneous receptions of inputs from both user and assistant. Inspired by Wang et al. (2025), OmniDRCA-Duplex organizes inputs into a double-channel format, specifically, an alternating sequence of user speech and assistant speech along the temporal axis. Besides, same position id is used in one time slot, which paired user and assistant speech representations together. The model enforces attention masking between concurrent speech representations in individual time slots since they appear simultaneously. This structured arrangement, termed time-division multiplexing (TDM) input stream, is visually illustrated in Figure 2 TDM-Duplex. This approach allows the model to accept user speech when the assistant is generating speech, effectively utilizing the time slots that would otherwise be idle. The TDM input stream is designed to handle both user and assistant speech inputs, ensuring a seamless interaction experience.
Awareness of Interruptions.
As our early experiments show, the model may not be aware of user interruptions during the generation of assistant speech, since the model prefers to generate speech based on the content in the time slots from its own role (that is, Assistant Speech Token Stream). This can lead to a lack of responsiveness to user interruptions and a less engaging interaction experience. To address this issue, we implement a mechanism that enhances the model’s awareness to user interruptions during assistant speech generation. We achieve this goal by adding an additional head to predict the assistant’s speech generation state. The model is trained to predict whether the assistant is currently generating speech or not, based on the input speech tokens. This additional head provides the model with contextual information about the assistant’s speech generation state, enabling it to respond more effectively to user interruptions.
A.2 Experiments
Experimental Setup.
The original speech datasets using for OmniDRCA post-training are edited to simulate full-duplex conversation scenarios. Two types of datasets are constructed: 1) Non-interruption Dataset: Dialogue turns are structured into dual-channel audio. To better simulate human interactions, a response time of $0\sim 400$ ms is randomly inserted between dialogue turns. This dataset enables the model to learn when to initiate and terminate speech. 2) Interruption Multi-Turn Dialogue Dataset: The assistant’s responses in dialogue turns are randomly truncated, followed by inserting a new user input , before constructing dual-channel datasets. Through this data simulation approach, the model learns to handle incomplete interrupted semantics. In total, about 4K hours non-interruption dataset and 4K hours interruption multi-turn dialogue dataset are adopted for OmniDRCA-Duplex training. UltraEval-Audio Benchmark is used for evaluating the semantic ability, while multi-turn Llama Questions randomly pairing two of questions together is used for evaluating the interruption ablitily. Specific, providing these two rounds Llama Questions examples, if the model can mute itself while the user is speaking and respond after the user finishes, it is considered a successful Turn-taking. The testing evaluates the success rate of Turn-taking. 666 Here, Turn-taking@k Metrics is not conducted unlike other duplex speech models (Zhang et al., 2024a). This approach stems from intentional dataset construction aimed at endowing the model with human-like response timing and enabling scenarios where semantic information coexists simultaneously in both channels. The design specifically emphasizes making the model’s turn-taking response time more flexible rather than prioritizing the speed of turn-taking as the objective for the duplex model.
Table 7: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks Comparing the original OmniDRCA model and its Duplex Model.
Model Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Moshi 7B 24 23.8 43.7 16.7 27.1
OmniDRCA-Duplex-Small 2B 33 16.1 46.6 13.6 27.3
OmniDRCA-Small 2B 38 22.1 55.3 17.9 33.3
Results.
As demonstrated in Table 7, evaluation results on the UltraEval-Audio Spoken Question Answering Benchmark indicate that OmniDRCA-Duplex-Small attains an average score of 27.3. This represents a 18.2% relative performance degradation compared to the baseline OmniDRCA-Small model’s average score of 33.3. Notably, the experimental results show that OmniDRCA-Small-Duplex (2B model size) has a slight performance advantage over the larger 7B full-duplex model Moshi (27.3 versus 27.1). For the model’s interruption performance, by setting the multi-turn Llama Questions with intervals of $t=3s$ between turns, OmniDRCA-Small-Duplex achieves 98.64% success rate of Turn-taking, meaning the model can keep silent when user is talking and consistently respond to user-interrupted utterances in the second round.
Table 8: The Impact of SRM Pretraining Data on Llama Questions Performance.
Model S2M (T/S) S2T T2M (T/S) T2T
OmniDRCA-Small (Less) 52.67 / 44.33 60.33 67.67 / 53.00 75.00
   w/o. SRM-PT 38.33 / 30.33 56.00 59.33 / 46.33 73.33
Figure 3: Performance Scaling of OmniDRCA-Small (w/o. Cont) on LLaMA Question Benchmark.
Appendix B Dataset Scaling
Pre-Training
To examine the impact of pretraining data on model performance, the SRM pretraining step is eliminated, and the OmniDRCA-Small model is retrained. For computational efficiency, fewer post-training steps are adopted, called OmniDRCA-Small (Less), leading to the inferior performance compared to the results of OmniDRCA-Small reported in the main text. As shown in Table 8, SRM pretraining has the most significant impact on the performance of S2M and T2M, while its effect on S2T and T2T is relatively minor. This is reasonable because SRM is primarily responsible for refining speech token generation tasks, thus exerting a smaller influence on purely textual generation. This underscores the importance of SRM pretraining.
Post-Training
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on OmniDRCA-Small (w/o. Cont). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
Appendix C Prompt Templates
System prompts for multimodal interaction patterns.
As shown in Table 9, there are five types of system prompts categorized based on the model’s output partitioning. During inference, specifying these system prompts enables the generation of desired output results.
Table 9: System Prompts for Multimodal Interaction.
Pattern Abbr. System Prompts
S2M & T2M You are a helpful assistant and asked to generate both text and speech tokens at the same time.
S2T & T2T You are a helpful assistant and asked to generate text tokens.
STC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SAC You are a helpful assistant. Let’s think step by step. Think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SUC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, and then think of both appropriate text and speech responses at the same time.
Evaluation Templates.
To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, datasets from two established evaluation frameworks were adopted: UltraEval-Audio [^7] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench [^8] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. For UltraEval-Audio, string-matching algorithms were applied to Llama Questions, Web Questions, and TriviaQA, while AlpacaEval utilized GPT-4o-mini with the Alpaca scoring template. In OpenAudioBench, GPT-4o-2024-08-06 was employed for scoring with different templates shown in Figures, all in strict compliance with the tools’ prescribed methodologies.
AlpacaEval Evaluation Template [Instruction] Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]". [Question] (Instruction Transcription) [The Start of Assistant’s Answer] (Response Transcription) [The End of Assistant’s Answer]
TriviaQA and Web Questions Evaluation Template Your will be given a question, the reference answers to that question, and an answer to be judged. Your tasks is to judge whether the answer to be judged is correct, given the question and reference answers. An answer considered correct expresses or contains the same meaning as at least **one of** the reference answers. The format and the tone of the response does not matter. You should respond in JSON format. First provide a one-sentence concise analysis for the judgement in field ‘analysis’, then your judgment in field ‘judgment’. For example, “‘json "analysis": "<a one-sentence concise analysis for the judgement>", "judgment": < your final judgment, "correct" or "incorrect"> ”’ # Question (instruction) # Reference Answer (targets) # Answer To Be Judged (answer_to_be_judged)
Llama Questions Evaluation Template ## Background You are a professional QA evaluation expert. You need to assess whether the model’s answer is correct based on the standard answer. ## Scoring Criteria Correct: The answer matches or is equivalent to the standard answer Incorrect: The answer is wrong or irrelevant to the question ## Evaluation Guidelines 1. The expression of answers can be flexible, not requiring exact matches. For example: - Numbers can be expressed in either Arabic numerals or words - Proper nouns can be in either English or Chinese - Differences in punctuation can be ignored 2. Focus on whether the core meaning of the answer is correct ## Output Format Provide the reasoning for your score, then generate the result in "[]" format and make sure it contains "the score is [Correct]" or "the score is [Incorrect]", for example: “‘ The answer is correct and equivalent to the standard answer, the score is [Correct] ”’ or “‘ The answer is incorrect and does not match the standard answer, the score is [Incorrect] ”’ ## Question: (instruction) ## Standard Answer: (targets) ## Model’s Answer: (answer_to_be_judged)

### v2

@@
-OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment
+DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations
 Abstract
@@
 (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation.
-This paper presents OmniDRCA, a parallel speech-text foundation model based on joint autoregressive modeling, featuring dual-resolution speech representations and contrastive cross-modal alignment.
-Our approach processes speech and text representations in parallel while enhancing audio comprehension through contrastive alignment.
-Experimental results on Spoken Question Answering benchmarks demonstrate that OmniDRCA establishes new state-of-the-art (SOTA) performance among parallel joint speech-text modeling based foundation models, and achieves competitive performance compared to interleaved models.
-Additionally, we explore the potential of extending the framework to full-duplex conversational scenarios.
-[^1] Our implementation will be released at https://github.com/lxchtan/OmniDRCA.
+This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations.
+Whereas current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz.
+Experimental results on Spoken Question Answering benchmarks demonstrate that DrVoice establishes new state-of-the-art (SOTA) performance among similar size speech foundation models with relative small amount of data.
 1 Introduction
@@
 E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
-Recent research efforts on E2E models to address the above challenges have focused on two primary directions: Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models.
-Text-Driven Speech Models refer to systems in which LLMs ingest speech representations as input, produce textual responses, and simultaneously utilize the LLM’s latent representations to synthesize audio.
-In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and autoregressively generating both text tokens and speech tokens.
-The key distinction lies in whether the generated speech tokens are fed back autoregressively into the LLM during the generation process.
-Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024).
-Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM, ensuring temporal alignment.
+Recent research efforts on E2E models to address the above challenges have focused on two primary directions Chen et al.
+(2024): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models.
+Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation.
+In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously.
+The key distinction lies in whether the generation of speech tokens can influence the subsequent generation of text tokens within the LLM.
+This feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models.
+Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024; KimiTeam et al., 2025).
+Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM.
 While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations.
-Text-Driven Speech Models prevent LLMs from perceiving generated speech tokens during inference, thereby limiting dynamic adjustment of text generation.
+The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins.
+This prevents the LLM from conditioning its textual output on the generated acoustic tokens, thus limiting its ability to control fine-grained paralinguistic attributes like emotion and prosody.
 Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge.
-Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials; hence, in this work, we focus on enhancing joint speech-text modeling.
-In this work, we propose OmniDRCA, a novel parallel speech-text foundation model.
-Our architecture preserves end-to-end advantages while significantly improving speech-text semantic alignment through two key innovations: 1) Dual-Resolution Speech Representations (DRSR): DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens.
-For the LLM’s speech comprehension, we introduce a grouping mechanism, mapping 25Hz discrete audio tokens to 5Hz speech representations that align much better with textual granularity, thereby leveraging LLM’s capabilities more effectively and improving speech understanding.
-Combined speech-text embeddings form the assistant’s autoregressive input.
-The multimodal LLM (MLLM) predicts text tokens.
-For speech generation, latent representations are ungrouped and autoregressively processed by the speech refinement model (SRM) for individual speech token prediction.
-During inference, due to input-output resolution discrepancies, each MLLM inference triggers $k$ parallel SRM executions, where $k$ is the grouping factor.
-2) Contrastive Cross-modal Alignment (CCA): Similarity between grouped speech representations and text representations is computed to optimize their mutual distances.
-While grouping aligns speech-text granularity, CCA further enhances semantic consistency between modalities before they enter the LLM and further enhances speech comprehension.
-Considering inference efficiency for varying output requirements in practical applications, we also design system prompts to control the output modes, including text-only output, parallel speech-text output, and parallel speech-text output after chain-of-modality (CoM) (Zhang et al., 2023a).
-Experimental results demonstrate the efficacy of CoM in improving performance.
+Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials Chen et al.
+(2024); hence, in this work, we focus on enhancing joint speech-text modeling.
+Recently, Kimi-Audio KimiTeam et al.
+(2025) has set a new state-of-the-art (SOTA) in joint speech-text modeling.
+However, this approach faces notable challenges.
+It not only requires extensive data for training but also incurs significant computational costs due to its 12.5Hz audio representation.
+Furthermore, this high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
+In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR), which compresses the input down to 5Hz.
+DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens.
+For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, which better match the pace of text.
+During generation, combined speech-text embeddings form the assistant’s autoregressive input.
+The hidden states captured from shared LLM layers are then passed in parallel to a text head for text token prediction, and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens.
+Due to input-output resolution discrepancies, each LLM inference triggers $k$ parallel SRH executions, where $k$ is the grouping factor.
+To further enhance the model’s reasoning and the coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy.
+This approach prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation.
+This intermediate reasoning step serves to improve modality alignment and reduce errors.
+For practical applications, we design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output.
+Experimental results demonstrate the efficacy of CoM in significantly improving performance.
 For reproducibility, we will release all source code and model checkpoints.
-Our contributions are three-fold: 1) We propose OmniDRCA, a novel parallel speech-text foundation model that enhances speech-to-speech generation based on discrete speech tokens and LLMs, through dual-resolution speech representations and contrastive cross-modal alignment.
-2) Comprehensive experiments demonstrate OmniDRCA’s superiority over the state-of-the-art (SOTA) parallel model LUCY (Gao et al., 2025) and competitive performance against SOTA interleaved models across multiple spoken question answering benchmarks.
-3) We further explore the potential of OmniDRCA for full-duplex voice conversation.
-By leveraging time-division multiplexing, OmniDRCA-Duplex supports real-time responses to user interruptions with minimal performance degradation.
+Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR).
+This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities.
+2) We introduce a CoM-Mixing training strategy which acts as a curriculum, using the structured reasoning of CoM to scaffold directly speech generation task.
+3) Through comprehensive experiments, we demonstrate that DrVoice achieves new state-of-the-art (SOTA) performance on multiple spoken question answering benchmarks, validating the effectiveness of our proposed architecture and strategies.
 2 Related Work
@@
 While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023).
-In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to its competitive semantic capabilities and support for streamlined streaming speech generation.
+In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to their competitive semantic capabilities and compatibility with a more lightweight, diffusion-free vocoder for direct speech synthesis.
 E2E Speech Foundation Models.
@@
 Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech.
-The contemporaneous Text-Driven model Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production.
+Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production.
 Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
 Joint Speech-Text Models explore two different architectures, Interleaved and Parallel.
-Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024) conduct parallel decoding.
+Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024; KimiTeam et al., 2025) conduct parallel decoding.
+The interleaved model VITA-Audio Long et al.
+(2025) achieves efficient audio generation by introducing multiple cross-modal token prediction (MCTP) modules.
+It demonstrates superior performance across multiple benchmarks while maintaining high-quality speech synthesis.
 The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations.
-Our approach differs from Moshi with two key distinctions: (1) the $k$ tokens correspond to grouped semantic speech tokens rather than disparate acoustic tokens, and (2) our speech refinement module explicitly ingests full-sequence contextual information to enhance semantic coherence.
-While SLAM-Omni (Chen et al., 2024) proposes grouped semantic token prediction for speech generation, our approach designs an ungrouping mechanism and incorporates a Speech Refined Model for autoregressive generation of individual speech tokens.
-Another contemporaneous parallel model Kimi-Audio (KimiTeam et al., 2025) is closest to our approach.
-To mitigate limitations of discrete tokens in audio understanding, they introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper features, preserving both semantic and acoustic information.
-In contrast, our approach enhances discrete token comprehension through contrastive loss and grouping strategies, thereby improving LLM’s understanding of speech tokens while maintaining representational efficiency.
-Contrastive Cross-Modal Alignment.
-CLIP (Radford et al., 2021) aligns image and text representations through contrastive pretraining.
-VILA-U (Wu et al., 2024) integrates contrastive loss and reconstruction loss during vision token pretraining to enhance feature alignment.
-In speech research, Ye et al.
-(2022) leverages multi-class N-pair contrastive loss (Sohn, 2016) for speech translation.
-Zheng et al.
-(2024) leverages InfoNCE loss (He et al., 2020) for spoken language understanding.
-We incorporate contrastive loss to strengthen speech-text semantic alignment.
-Crucially, we adopt gradient stop (Chen and He, 2021) on text embeddings during contrastive loss computation – this modification preserves speech representation learning while preventing text feature distortion.
+To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper Radford et al.
+(2023) features, preserving both semantic and acoustic information.
+In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
 3 Methodology
-Figure 1 illustrates the architecture of OmniDRCA.
-The system consists of two main components: (1) a Multimodal Large Language Model (MLLM) designed to process interleaved speech-text inputs and achieve cross-modal alignment, and (2) a Speech Refined Model (SRM) specialized in high-fidelity speech synthesis.
-The system operates through multimodal input encoding, joint representation learning with contrastive alignment, and coordinated speech-text output generation.
-During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce multimodal context embeddings, and decoded into synchronized speech-text responses through SRM and the language model head.
+Figure 1 illustrates the architecture of DrVoice.
+The system consists of three main components: (1) Speech Tokenizer and Speech Encoder are designed to process the wave of speech into hidden representation, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, text head, and speech refined head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens.
+The system operates through multimodal input encoding and coordinated speech-text output generation.
+During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head.
+To effectively train this system, we introduce the CoM-Mixing Training strategy.
 3.1 Speech Tokenization and Detokenization
@@
 For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency.
-Figure 1: Overview of OmniDRCA.
-User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text token prediction.
-Hidden representations are ungrouped and processed by the SRM through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.
+Inspired by Kimi-Audio KimiTeam et al.
+(2025) and aiming for improved audio understanding, we utilize the Whisper-Large-v3 Speech Encoder to extract continuous audio representations.
+Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM.
+Finally, the processed continuous representations are stacked with discrete ones and fed into the LLM.
+Figure 1: Overview of DrVoice.
+User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction.
 The generated speech tokens are then converted to speech waveform by the speech detokenizer.
-We omit speech tokenizer and detokenizer for simplicity.
+Note that SRH through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.
+Table 1: Multimodal Interaction Patterns.
+Pattern Name Abbr. Modality Flow
+Speech-to-Multimodal S2M Speech $\rightarrow$ Joint speech-text response
+Speech-to-Text S2T Speech $\rightarrow$ Text-only response
+Text-to-Multimodal T2M Text $\rightarrow$ Joint speech-text response
+Text-to-Text T2T Text $\rightarrow$ Text-only response
+Speech-Text Chain STC Speech $\rightarrow$ Text transcription $\rightarrow$ Text response $\rightarrow$ Multimodal response
+Speech-Assisted Chain SAC Speech $\rightarrow$ Text response (agent perspective) $\rightarrow$ Multimodal response
+Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
 3.2 Multimodal Large Language Model (MLLM)
@@
 Parallel Joint Speech-Text Model.
-Leveraging the autoregressive generation capability of MLLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the MLLM at each timestep.
+Leveraging the autoregressive generation capability of LLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the shared LLM layer at each timestep.
 Their embeddings are added as model input, forming a parallel speech-text architecture.
@@
 $c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
-where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. The autoregressive generation process is as follows:
+where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively.
+Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with special token <|SIL|>to align them for each utterance.
+The autoregressive generation process is as follows:
 $P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
@@
 $\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
-[…]
-Cross-modal Contrastive Alignment.
-To enable cross-modal alignment, a contrastive learning objective aligns grouped speech representations with their textual counterparts by maximizing their cosine similarity and computing a contrastive loss.
-This joint optimization ensures coordinated positioning of both modalities within the shared representation space.
-$\mathcal{L}_{\text{cont}}=-\cos(E_{\text{text}}(\mathbf{t}),E_{\text{speech}}(\mathbf{g})),$ (4)
-where $E_{\text{text}}$ and $E_{\text{speech}}$ are the mean of embeddings of text tokens and grouped speech tokens between sample pairs, respectively.
-Different from early applications of contrastive alignment between speech and text (Zheng et al., 2024), our approach eliminates negative sample computation and prevents mode collapse through gradient stop on the text branch.
-3.3 Speech Refined Model (SRM)
-SRM is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the MLLM as conditional input and incorporates contextual speech information to autoregressively generate speech tokens.
+where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates.
+This design compresses the speech token length from $T$ to $T/k$ and the grouped speech representations align better with text.
+Different from Chen et al.
+(2024) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
+Speech Refined Head.
+SRH is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens.
 While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details.
-To address this limitation, an ungrouping process is performed as follows: the MLLM’s final hidden state is mapped to group-sized embeddings via linear projection
-$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[MLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (5)
+To address this limitation, an ungrouping process is performed as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
+$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[SLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (4)
 and time splitting
-$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}]\quad\text{where}\quad\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k},$ (6)
-the resulting $\mathbf{H}$ serves as conditional input for SRM that autoregressively generates speech tokens.
-Following our practice on MLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated.
-SRM is trained to maximize the conditional probability:
-$\mathcal{L}_{\text{SRM}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}),$ (7)
+$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}],$ (5)
+where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$, the resulting $\mathbf{H}$ serves as conditional input for SRH that autoregressively generates speech tokens.
+Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated.
+Speech token prediction is trained to maximize the conditional probability:
+$\mathcal{L}_{\text{SRH}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}_{<i}),$ (6)
 where $s_{i}$ represents the $i$-th speech token.
-By optimizing this objective, SRM learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from MLLM.
-This design enables SRM to effectively leverage the semantic and acoustic information encoded in the hidden representations of MLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
-The end-to-end training objective integrates the three losses through multi-task learning paradigm:
-$\mathcal{L}_{\text{total}}=\lambda\mathcal{L}_{\text{LM}}+\mu\mathcal{L}_{\text{cont}}+\gamma\mathcal{L}_{\text{SRM}},$ (8)
-$\mathcal{L}_{\text{LM}}=-\sum_{i=1}^{T}\log P(t_{i}|t_{<i},\mathbf{g}),$ (9)
-where $\mathcal{L}_{\text{LM}}$ is the autoregressive loss over MLLM, $\lambda$, $\mu$, and $\gamma$ are hyperparameters.
-3.4 Multimodal Interaction Patterns
-Table 1: Multimodal Interaction Patterns.
-Pattern Name Abbr. Modality Flow
-Speech-to-Multimodal S2M Speech $\rightarrow$ Joint speech-text response
-Speech-to-Text S2T Speech $\rightarrow$ Text-only response
-Text-to-Multimodal T2M Text $\rightarrow$ Joint speech-text response
-Text-to-Text T2T Text $\rightarrow$ Text-only response
-Speech-Text Chain STC Speech $\rightarrow$ Text transcription $\rightarrow$ Text response $\rightarrow$ Multimodal response
-Speech-Assisted Chain SAC Speech $\rightarrow$ Text response (agent perspective) $\rightarrow$ Multimodal response
-Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
-Practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs.
-Furthermore, Zhang et al.
-(2023a) found that the chain-of-modality (CoM) methodology can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
+By optimizing this objective, SRH learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from SLLM.
+This design enables SRH to effectively leverage the semantic and acoustic information encoded in the hidden representations of SLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
+The end-to-end training objective integrates the two losses through multi-task learning paradigm:
+$\mathcal{L}_{\text{MLLM}}=\lambda\mathcal{L}_{\text{TH}}+\mu\mathcal{L}_{\text{SRH}},$ (7)
+$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
+where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over text head, $\lambda$ and $\mu$ are hyperparameters.
+3.3 CoM-Mixing Training strategy
+The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
+Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs.
 Seven interaction patterns for diverse input-output requirements are summarized in Table 1.
@@
 For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs.
-Detailed prompts are shown in the Appendix.
-For training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model.
+Detailed prompts are shown in the Appendix A.
+During inference, specifying these system prompts enables the generation of desired output results.
+For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model.
 During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement.
@@
 4 Experiments
-Table 2: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks.
-Baseline results are from Yao et al.
-(2024) except LUCY (S2) and GLM-4-Voice-Small from Gao et al.
-(2025) and Zeng et al.
-(2025) respectively.
+Table 2: Speech-to-Text (S $\rightarrow$ T) and Speech-to-Speech (S $\rightarrow$ S) generation performance on spokenQA benchmarks.
+The results for the S2T baselines are taken from the previous SOTA model Kimi-Audio’s paper (Xu et al., 2025), and the S2S results are taken from the previous SOTA model VITA-Audio’s paper (Long et al., 2025).
+Otherwise, results are taken from their original papers.
+The model size refers only to the parameter count of the LLM backbone, excluding the audio encoder, audio tokenzier, and audio detokenzier module.
+Frame Rate (In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone.
+The average number of text tokens corresponding to one second of speech is denoted by $\tau$.
 Metrics are benchmark-specific (higher is better).
 The Avg.
-score averages four benchmarks (Gray: AlpacaEval excluded).
-Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg.
-GPT-4o-Realtime Closed-source 74 51.6 71.7 69.7 66.7
-Step-Audio-Chat Text-Driven 130B 43 56.1 56.7 36.7 48.1
-MiniCPM-o 2.6 Text-Driven 8B 51 40.0 61.0 40.2 48.1
-Llama-Omni Text-Driven 8B 39 22.9 45.3 10.7 29.5
-VITA-1.5 Text-Driven 7B 20 28.1 46.7 23.3 29.5
-GLM-4-Voice Interleaved 9B 51 32.0 50.0 36.4 42.3
-GLM-4-Voice-Small Interleaved 1.5B 5.4 18.3 4.6 9.4
-LUCY (S2) Parallel 7B 25.6 54.0 22.8 34.1
-Moshi Parallel 7B 24 23.8 43.7 16.7 27.1
-Mini-Omni Parallel 0.5B 25 12.8 22.0 6.9 16.7
-OmniDRCA Parallel 8B 45 30.0 65.0 32.9 43.2
-OmniDRCA-Small Parallel 2B 38 22.1 55.3 17.9 33.3
-Table 3: Speech-to-Text ($S\rightarrow T$) Generation Performance on OpenAudioBench using benchmark-specific metrics (higher is better).
-Baseline results are from KimiTeam et al.
-(2025), except LUCY (S2) and Moshi from Gao et al.
-(2025) and MiniCPM-o 2.6 from our experiments.
-Reasoning QA is removed since our model is only trained on English data.
-Avg.
-is averaged over the four benchmarks (Gray: AlpacaEval excluded).
-Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
-Qwen2-Audio-Chat Cascaded 7B 57.2 45.2 69.7 40.3 53.1
-MiniCPM-o 2.6 Text-Driven 8B 64.1 69.2 78.0 63.0 68.5
-Step-Audio-chat Text-Driven 130B 56.5 73.0 72.3 56.8 64.7
-Baichuan-Omni-1.5 Interleaved 7B 59.6 58.7 74.3 55.4 62.0
-GLM-4-Voice Interleaved 9B 57.9 55.4 76.0 51.8 60.3
-Moshi Parallel 7B 26.6 62.3 22.8 37.2
-LUCY (S2) Parallel 7B 29.3 57.3 23.3 36.6
-OmniDRCA Parallel 8B 58.0 51.7 79.7 47.7 59.7
-Table 4: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following (Zeng et al., 2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text.
-Baseline scores are copied from (Zeng et al., 2025), except MiniCPM-o 2.6 results are from our experiments.
+score averages two benchmarks.
+The best results in each group are highlighted in bold, while the second-best results are underlined.
+Model Size Frame Rate (In/Out) Llama Q. Web Q. Avg.
+$\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$
 […]
-MiniCPM-o 2.6 Text-Driven 8B 4.18 13.31
-Llama-Omni Text-Driven 8B 3.92 9.18
-GLM-4-Voice Interleaved 9B 4.33 7.83
-Moshi Parallel 7B 3.90 7.95
-Mini-Omni Parallel 0.5B 3.17 25.28
-OmniDRCA Parallel 8B 3.86 6.75
-OmniDRCA-Small Parallel 2B 3.87 8.86
-Table 5: Ablation study of Dual-Resolution Grouping Strategy Conducted on OmniDCRA-Small with Llama Questions Dataset.
-Modal Group Size S2T T2M (T/S) T2T
-S2T+T2T (Understanding) 1 55.33 61.33
-5 62.33 69.67
-T2M+T2T (Generation) 1 51.67 / 41.00 58.33
-5 42.00 / 25.33 64.00
-Table 6: Ablation study of Contrastive Cross-modal Alignment (Cont), SRM, and Multimodal Architecture Design (MD) on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
+DrVoice-Small (Our) 2B 5/5 73.0 56.0 42.1 31.1 57.5 43.5
+Moshi (2024) 7B 12.5/12.5 62.3 21.0 26.6 9.2 44.5 15.1
+[…]
+[…]
+[…]
+[…]
+[…]
+LUCY (S2) (2025) 7B 12.5/48 57.3 51.0 29.3 18.2 43.3 34.6
+[…]
+Kimi-Audio (2025) 8B 12.5/12.5 79.3 - 70.0 - 74.7 -
+[…]
+[…]
+Stream-Omni (2025) 7B 25/25 76.3 65.0 44.2 27.5 60.3 46.3
+[…]
+DrVoice (Our) 8B 5/5 85.0 71.0 65.2 40.1 75.1 55.5
+Table 3: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following Zeng et al.
+(2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text.
+Baseline scores are copied from Zeng et al.
+(2025).
+FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech.
+The best results in each group are bolded.
+[…]
+Mini-Omni (2024a) 0.5B 50/48 3.17 25.28
+DrVoice-Small 2B 5/5 3.87 8.86
+[…]
+Moshi (2025) 7B 12.5/12.5 3.90 7.95
+[…]
+DrVoice 8B 5/5 3.86 6.75
+Table 4: Impact of the Grouping Factor on Llama Q.
+S2T and S2M performance.
+Experiments were conducted on DrVoice-Small w/o.
+CSE, using data subsets for faster exploration.
+The S2T model was trained exclusively on mixed S2T and T2T data, while the S2M model was trained on mixed S2M and T2T data.
+Grouping Factor S2T S2M (T/S)
+1 55.67 4.00 / 2.67
+3 64.67 15.67 / 5.00
+5 63.33 37.67 / 28.00
+7 62.67 36.00 / 16.67
+Table 5: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
 Model S2M (T/S) S2T T2M (T/S) T2T STC (T/S) SAC (T/S) SUC (T/S)
-OmniDRCA-Small 64.33 / 55.33 64.67 69.67 / 58.67 69.67 67.33 / 59.67 65.67 / 57.00 67.67 / 57.00
-w/o Cont 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
-   w/o SRM 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
-   w/o MD 58.00 / 49.00 58.00 69.33 / 55.00 68.33
+DrVoice-Small 68.67 / 56.00 72.33 72.33 / 56.00 75.33 75.67 / 68.33 71.67 / 62.67 73.33 / 62.00
+w/o. CSE 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
+   w/o. SRH-Pretraining 38.33 / 30.33 56.00 59.33 / 46.33 73.33 67.33 / 57.67 54.00 / 42.33 54.33 / 42.67
+    w/o. SRH 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
+   w/o. CoM-Mixing 58.00 / 49.00 58.00 69.33 / 55.00 68.33 –/– –/– –/–
+Figure 2: Computational Resources across different Grouping Factor.
+Figure 3: Performance Scaling of DrVoice-Small (w/o. Continuous Speech Encoder) on LLaMA Question Benchmark.
 4.1 Experimental Setup
 Datasets.
-Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for SRM pre-training.
-17.8K hours of English speech conversations data synthesized from text are adopted for OmniDRCA training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^2] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b).
-Following prior works (Yao et al., 2024; Li et al., 2025), we evaluate the performance on the widely used Spoken Question Answering benchmarks, including AlpacaEval (Li et al., 2023), Web Questions (Berant et al., 2013), Llama Questions (Nachmani et al., 2024), and TriviaQA (Joshi et al., 2017).
-To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, we adopt datasets from two established evaluation frameworks: UltraEval-Audio  [^3] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench  [^4] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation.
-As discussed in Huang et al.
-(2025); Chen et al.
-(2025), the results on TriviaQA are for reference only and marked with “*” due to the absence of an official test set.
-Moreover, we detect performance anomalies of OmniDRCA on the UltraEval-Audio TriviaQA dataset; therefore, we replace UltraEval-Audio TriviaQA dataset with OpenAudioBench TriviaQA dataset for a more meaningful evaluation  [^5] For reference, this setup yields a score of 42.6 for MiniCPM-o 2.6 while its score on the original UltraEval-Audio TriviaQA is 40.2..
+Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for pre-training speech refined head.
+17.8K hours of English speech conversations data synthesized from text are adopted for DrVoice training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^1] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b).
+Following the common practice in prior works (Long et al., 2025; KimiTeam et al., 2025; Yao et al., 2024; Zhang et al., 2025), the Speech-to-Speech ($S\rightarrow S$) and Speech-To-Text ($S\rightarrow T$) performance is evaluated on widely used Spoken Question Answering benchmarks, including Web Questions (Berant et al., 2013) and Llama Questions (Nachmani et al., 2024).
 Evaluation Metrics.
-Following the settings of UltraEval-Audio and OpenAudioBench, we utilize G-Eval (Liu et al., 2023) to evaluate model performance on AlpacaEval, and employ accuracy to assess model performance on other datasets.
+Accuracy is emplyed to assess model performance on datasets.
 The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al.
@@
 Baselines.
-Representive baselines are selected to cover diverse modeling paradigms: the closed-source model GPT-4o-Realtime (OpenAI, 2024b), cascaded system Qwen2-Audio-Chat (Chu et al., 2024), Text-Driven models including Step-Audio-Chat (Huang et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni (Fang et al., 2024), and VITA-1.5 (Fu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025) and GLM-4-Voice (Zeng et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Mini-Omni (Xie and Wu, 2024a).
+Representive baselines are selected to cover diverse modeling paradigms: the cascaded system Audio Flamingo 3 (Zhang et al., 2025), Text-Driven models including Freeze-Omni (Défossez et al., 2024), MinMo (Chen et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni2 (Fang et al., 2025), and Qwen2.5-Omni (Xu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025), GLM-4-Voice (Zeng et al., 2025), Stream-Omni (Zhang et al., 2025), and VITA-Audio (Long et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Kimi-Audio (Xu et al., 2025).
+Noting that Kimi-Audio achieves state-of-the-art (SOTA) performance on Speech-to-Text Question Answering, while VITA-Audio sets the SOTA on Speech-to-Speech Question Answering.
 This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
 Implementation Details.
-SRM uses Qwen2.5-0.5B as its base LLM, trained following $T2M$ paradigm.
-OmniDRCA uses Qwen2.5-7B-Instruct as its base LLM, while OmniDRCA-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM.
+Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT).
+DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM.
 Grouping factor is set to $k=5$.
 The warmup rate is set to $1\%$ of the total training steps.
-For OmniDRCA, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for OmniDRCA-Small and SRM it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$.
+For DrVoice, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for DrVoice-Small and SRH-PT it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$.
 The AdamW method (Loshchilov and Hutter, 2019) is used for optimization.
-Experiments are run on 32x NVIDIA Tesla A100 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding.
-For OmniDRCA, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory.
-It takes about $20$ hours on SRM pretraining, and about $20$ hours on OmniDRCA post training.
+Experiments are run on 32x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding.
+For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory.
+It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post training.
 4.2 Results and Discussions
-Tables 2 and 3 compare OmniDRCA with the representative and competitive baselines across all major model categories, on speech-to-speech and speech-to-text generation performance.
-Table 2 shows that in prior works, best Text-Driven models substantially outperform top-tier Joint Speech-Text models, because text-driven systems do not allow speech tokens to interfere with the generation process, thereby minimally impacting the LLM’s generation paradigm.
-Moreover, prior interleaved models often outperform parallel models, though they exhibit higher latency (e.g., 13 text tokens for first speech in GLM-4-Voice) and lack temporal alignment, complicating multistream duplex modeling.
-As shown in Tables 2 and 3, OmniDRCA achieves SOTA performance under the parallel paradigm, achieving a substantial 24.9% relative gain over LUCY (S2) on the average score across three Speech-to-Speech QA benchmarks and 60.4% gain over Moshi across three Speech-to-Text QA benchmarks, excluding AlpacaEval in both cases.
-OmniDRCA also shows competitive performance against the interleaved GLM-4-Voice and Baichuan-Omni-1.5.
-However, a relatively large gap still remains when compared to text-driven systems.
-In future work, we plan to train the model on more extensive datasets to improve the performance.
-The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (WER(%)), as shown in Table 4.
-Among models with comparable sizes, text-driven models tend to exhibit higher WER, while joint speech-text models achieve relatively lower WER.
-The model size also significantly impacts WER.
-OmniDRCA achieves the lowest WER, indicating strong alignment when generating both speech and text simultaneously.
+[…]
+The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (ASR-WER(%)), as shown in Table 3.
+Among models with comparable sizes, DrVoice achieves low ASR-WER, indicating strong alignment when generating both speech and text simultaneously.
 In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a).
-The gap in UTMOS between OmniDRCA and GLM-4-Voice suggests room for improvement in modeling prosody and preserving acoustic details.
+The gap in UTMOS between DrVoice and GLM-4-Voice suggests room for improvement in speech naturalness and intelligibility.
 We conduct extensive ablation study to analyze effects of critical algorithmic designs.
-Dual-Resolution Grouping Strategy.
-Table 5 reveals the dual effects of our grouping strategy: while improving speech understanding (12.6% relative gain in S2T with group size 5) and text capabilities (13.6% relative gain in T2T), it degrades speech generation quality by 38.2% relative in T2M.
-On the other hand, as shown in Table 6, SRM demonstrates remarkable effectiveness in speech token generation, achieving 245.9% and 71.4% relative improvements in S2M and T2M tasks respectively.
-Notably, SRM also enhances speech comprehension metrics (S2T and STC-T) by 10-16% relative, suggesting that improved speech token reconstruction capability improves latent speech representations.
-The text-only performance (T2T) remains stable, confirming SRM’s modality-specific effectiveness without interfering with text processing pathways.
-Our dual-resolution architecture resolves this conflict by applying grouping only in the comprehension phase (5Hz) while using raw frame resolution (25Hz) with SRM for generation.
-Contrastive Cross-modal Alignment.
-As shown in Table 6, the introduction of speech-text contrastive learning for sentence-level semantic alignment substantially enhances the model’s semantic understanding capabilities.
-While the final evaluation metric (S2S performance) shows improvement over the counterpart without contrastive loss, the enhancement exhibits modality-specific efficacy.
-For instance, S2T performance improves by 2.34% absolute (64.67 over 62.33), while T2T performance degrades by 4.33%, suggesting that while improving speech comprehension, the contrastive objective partially compromises the LLM’s native text understanding capacity.
-Our multimodal evaluation framework effectively reveals these nuanced impacts, guiding future work on mitigating text performance degradation through more balanced loss design.
-Multimodal Architecture Design (MD).
-Our multimodal design achieves two key goals: (1) context-aware modality selection and (2) implicit data augmentation.
-As shown in Table 6, on the speech generation task, STC/SAC/SUC demonstrate substantial improvements over S2M.
-STC achieves 61.00, substantially surpassing the direct generation baseline of 55.33.
-This reveals the critical role of CoM while also indicating potential optimization opportunities for our S2M model.
-Without incorporating the three CoM variants (STC/SAC/SUC), the model fails to adopt this generation paradigm guided by system prompts.
-Final S2M performance shows that training without MD results in a 7.5% performance degradation, confirming the data augmentation effect of our approach.
-Inspired by Wang et al.
-(2025), to enable full-duplex interaction, we propose a time-division multiplexing (TDM) input stream that allows OmniDRCA to accept user speech inputs during the assistant’s speech generation phase, thereby efficiently utilizing otherwise idle time slots.
-The resulting OmniDRCA-Duplex achieves full-duplex dialogue capabilities.
-The methodology and empirical validations of OmniDRCA-Duplex are detailed in Appendix.
-5 Conclusions, Limitations, and Future Work
-We introduce OmniDRCA, a novel parallel speech-text foundation model that integrates joint autoregressive modeling with dual-resolution speech representations and contrastive cross-modal alignment.
-Experiments show that OmniDRCA sets new SOTA speech-to-speech/text generation performance among parallel models and reaches competitive performance to interleaved models.
-Limitations: (1) OmniDRCA’s speech generation quality remains constrained by suboptimal speech tokenization.
-We will explore other tokenization methods, such as CosyVoice’s speech tokenizer, to address this limitation.
-(2) We will extend exploration from current speech understanding and generation tasks to broader audio comprehension/synthesis tasks, and integrate visual modality.
-Appendix A OmniDRCA-Duplex
-As mentioned in Introduction, we explore the potential of OmniDRCA for full-duplex voice conversation.
-By leveraging time-division multiplexing, we introduce a novel spoken dialogue model, OmniDRCA-Duplex, which supports real-time responses to user interruptions with minimal performance degradation.
-In this section, we first introduce OmniDRCA-Duplex in Appendix A.1, then present the experimental results and discussions of its full-duplex conversation capabilities in Appendix A.2.
-A.1 Methodology
-As shown in Figure 2, compared with OmniDRCA, the overall architecture of OmniDRCA-Duplex remains largely unchanged, which is also composed of the multimodal LLM (MLLM) and the speech refinement model (SRM).
-The primary differences between OmniDRCA-Duplex and the original OmniDRCA model lie in two aspects.
-The first aspect is the organization of the input structure.
-OmniDRCA-Duplex organizes user speech and assistant speech alternately along the temporal axis as input, referred to as the Time-Division Multiplexing Input Stream, while OmniDRCA adopts turn-based speech as the same as text-based LLM.
-For the second aspect, to enhance the model’s sensitivity to user interruptions, an additional assistant state prediction head is introduced in OmniDRCA-Duplex to determine whether muting should be activated.
-The following paragraphs provide detailed descriptions of these two designs.
-Figure 2: The architecture of OmniDRCA-Duplex.
-Time-Division Multiplexing Input Stream.
-Full-duplex spoken dialogue models allow simultaneous receptions of inputs from both user and assistant.
-Inspired by Wang et al.
-(2025), OmniDRCA-Duplex organizes inputs into a double-channel format, specifically, an alternating sequence of user speech and assistant speech along the temporal axis.
-Besides, same position id is used in one time slot, which paired user and assistant speech representations together.
-The model enforces attention masking between concurrent speech representations in individual time slots since they appear simultaneously.
-This structured arrangement, termed time-division multiplexing (TDM) input stream, is visually illustrated in Figure 2 TDM-Duplex.
-This approach allows the model to accept user speech when the assistant is generating speech, effectively utilizing the time slots that would otherwise be idle.
-The TDM input stream is designed to handle both user and assistant speech inputs, ensuring a seamless interaction experience.
-Awareness of Interruptions.
-As our early experiments show, the model may not be aware of user interruptions during the generation of assistant speech, since the model prefers to generate speech based on the content in the time slots from its own role (that is, Assistant Speech Token Stream).
-This can lead to a lack of responsiveness to user interruptions and a less engaging interaction experience.
-To address this issue, we implement a mechanism that enhances the model’s awareness to user interruptions during assistant speech generation.
-We achieve this goal by adding an additional head to predict the assistant’s speech generation state.
-The model is trained to predict whether the assistant is currently generating speech or not, based on the input speech tokens.
-This additional head provides the model with contextual information about the assistant’s speech generation state, enabling it to respond more effectively to user interruptions.
-A.2 Experiments
-Experimental Setup.
-The original speech datasets using for OmniDRCA post-training are edited to simulate full-duplex conversation scenarios.
-Two types of datasets are constructed: 1) Non-interruption Dataset: Dialogue turns are structured into dual-channel audio.
-To better simulate human interactions, a response time of $0\sim 400$ ms is randomly inserted between dialogue turns.
-This dataset enables the model to learn when to initiate and terminate speech.
-2) Interruption Multi-Turn Dialogue Dataset: The assistant’s responses in dialogue turns are randomly truncated, followed by inserting a new user input , before constructing dual-channel datasets.
-Through this data simulation approach, the model learns to handle incomplete interrupted semantics.
-In total, about 4K hours non-interruption dataset and 4K hours interruption multi-turn dialogue dataset are adopted for OmniDRCA-Duplex training.
-UltraEval-Audio Benchmark is used for evaluating the semantic ability, while multi-turn Llama Questions randomly pairing two of questions together is used for evaluating the interruption ablitily.
-Specific, providing these two rounds Llama Questions examples, if the model can mute itself while the user is speaking and respond after the user finishes, it is considered a successful Turn-taking.
-The testing evaluates the success rate of Turn-taking.
-666 Here, Turn-taking@k Metrics is not conducted unlike other duplex speech models (Zhang et al., 2024a).
-This approach stems from intentional dataset construction aimed at endowing the model with human-like response timing and enabling scenarios where semantic information coexists simultaneously in both channels.
-The design specifically emphasizes making the model’s turn-taking response time more flexible rather than prioritizing the speed of turn-taking as the objective for the duplex model.
-Table 7: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks Comparing the original OmniDRCA model and its Duplex Model.
-Model Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
-Moshi 7B 24 23.8 43.7 16.7 27.1
-OmniDRCA-Duplex-Small 2B 33 16.1 46.6 13.6 27.3
-OmniDRCA-Small 2B 38 22.1 55.3 17.9 33.3
-Results.
-As demonstrated in Table 7, evaluation results on the UltraEval-Audio Spoken Question Answering Benchmark indicate that OmniDRCA-Duplex-Small attains an average score of 27.3.
-This represents a 18.2% relative performance degradation compared to the baseline OmniDRCA-Small model’s average score of 33.3.
-Notably, the experimental results show that OmniDRCA-Small-Duplex (2B model size) has a slight performance advantage over the larger 7B full-duplex model Moshi (27.3 versus 27.1).
-For the model’s interruption performance, by setting the multi-turn Llama Questions with intervals of $t=3s$ between turns, OmniDRCA-Small-Duplex achieves 98.64% success rate of Turn-taking, meaning the model can keep silent when user is talking and consistently respond to user-interrupted utterances in the second round.
-Table 8: The Impact of SRM Pretraining Data on Llama Questions Performance.
-Model S2M (T/S) S2T T2M (T/S) T2T
-OmniDRCA-Small (Less) 52.67 / 44.33 60.33 67.67 / 53.00 75.00
-   w/o. SRM-PT 38.33 / 30.33 56.00 59.33 / 46.33 73.33
-Figure 3: Performance Scaling of OmniDRCA-Small (w/o. Cont) on LLaMA Question Benchmark.
-Appendix B Dataset Scaling
-Pre-Training
-To examine the impact of pretraining data on model performance, the SRM pretraining step is eliminated, and the OmniDRCA-Small model is retrained.
-For computational efficiency, fewer post-training steps are adopted, called OmniDRCA-Small (Less), leading to the inferior performance compared to the results of OmniDRCA-Small reported in the main text.
-As shown in Table 8, SRM pretraining has the most significant impact on the performance of S2M and T2M, while its effect on S2T and T2T is relatively minor.
-This is reasonable because SRM is primarily responsible for refining speech token generation tasks, thus exerting a smaller influence on purely textual generation.
-This underscores the importance of SRM pretraining.
-Post-Training
-To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on OmniDRCA-Small (w/o.
-Cont).
+Continuous Speech Encoder.
+As shown in Table 5, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs.
+Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation.
+Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively).
+In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00.
+This confirms the CSE’s effectiveness in processing and representing speech information.
+Dual-Resolution.
+Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation.
+Grouping Factor.
+Table 4 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency.
+Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M).
+For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain).
+The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5.
+Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, suggesting the efficiency of grouping machinism.
+Speech Refinement.
+As shown in Table 5, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks.
+By comparing the model with SRH (w/o.
+SRH-Pretraining) to one without (w/o.
+SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33).
+The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways.
+Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
+To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model.
+As shown in Table 5, removing pretraining (comparing w/o.
+CSE with w/o.
+SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%.
+The effect on S2T is smaller (10.2% drop), and it is negligible for T2T.
+This underscores the critical importance of SRH pretraining for refining speech token generation.
+CoM-Mixing Training Strategy.
+As shown in Table 5, where tasks guided by contextual system prompts (STC/SAC/SUC) demonstrate substantial improvements over direct S2M generation.
+For example, STC (T) achieves a score of 75.67, significantly surpassing the S2M (T) baseline of 68.67.
+This shows the model successfully learns to adopt the generation paradigm guided by system prompts.
+The data augmentation effect is confirmed by ablating the CoM-Mixing strategy entirely.
+Training without it (w/o.
+CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
+Data Scaling
+To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o.
+Continous Speech Encoder).
 Each S2S instance can be augmented into 7 distinct multimodal interaction patterns.
@@
 The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
-Appendix C Prompt Templates
-System prompts for multimodal interaction patterns.
+5 Conclusions and Future Work
+We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations.
+Experimental results demonstrate that DrVoice sets a new state-of-the-art (SOTA) benchmark in both speech-to-speech and speech-to-text generation tasks against current baselines.
+Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance.
+Our future work will address three key areas: (1) Enhancing Speech Generation Quality: To overcome the limitations of the current suboptimal tokenization and further improve speech fidelity, we plan to explore advanced speech tokenizers, such as the one utilized in CosyVoice.
+This will enable the model to capture finer acoustic details and generate more natural-sounding audio.
+(2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations.
+Inspired by recent advancements like Parrot (Wang et al., 2025), we will investigate the use of a time-division multiplexing (TDM) input stream.
+This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue.
+(3) Expanding to General Audio and Multimodality: Finally, we aim to extend our research beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds.
+The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.
+Appendix A Prompt Templates
 As shown in Table 6, there are five types of system prompts categorized based on the model’s output partitioning.
 During inference, specifying these system prompts enables the generation of desired output results.
-Table 9: System Prompts for Multimodal Interaction.
 Pattern Abbr. System Prompts
@@
 SUC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, and then think of both appropriate text and speech responses at the same time.
-Evaluation Templates.
-To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, datasets from two established evaluation frameworks were adopted: UltraEval-Audio [^7] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench [^8] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation.
-For UltraEval-Audio, string-matching algorithms were applied to Llama Questions, Web Questions, and TriviaQA, while AlpacaEval utilized GPT-4o-mini with the Alpaca scoring template.
-In OpenAudioBench, GPT-4o-2024-08-06 was employed for scoring with different templates shown in Figures, all in strict compliance with the tools’ prescribed methodologies.
-AlpacaEval Evaluation Template [Instruction] Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below.
-Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response.
-Begin your evaluation by providing a short explanation.
-Be as objective as possible.
-After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]".
-[Question] (Instruction Transcription) [The Start of Assistant’s Answer] (Response Transcription) [The End of Assistant’s Answer]
-TriviaQA and Web Questions Evaluation Template Your will be given a question, the reference answers to that question, and an answer to be judged.
-Your tasks is to judge whether the answer to be judged is correct, given the question and reference answers.
-An answer considered correct expresses or contains the same meaning as at least **one of** the reference answers.
-The format and the tone of the response does not matter.
-You should respond in JSON format.
-First provide a one-sentence concise analysis for the judgement in field ‘analysis’, then your judgment in field ‘judgment’.
-For example, “‘json "analysis": "<a one-sentence concise analysis for the judgement>", "judgment": < your final judgment, "correct" or "incorrect"> ”’ # Question (instruction) # Reference Answer (targets) # Answer To Be Judged (answer_to_be_judged)
-Llama Questions Evaluation Template ## Background You are a professional QA evaluation expert.
-You need to assess whether the model’s answer is correct based on the standard answer.
-## Scoring Criteria Correct: The answer matches or is equivalent to the standard answer Incorrect: The answer is wrong or irrelevant to the question ## Evaluation Guidelines 1.
-The expression of answers can be flexible, not requiring exact matches.
-For example: - Numbers can be expressed in either Arabic numerals or words - Proper nouns can be in either English or Chinese - Differences in punctuation can be ignored 2.
-Focus on whether the core meaning of the answer is correct ## Output Format Provide the reasoning for your score, then generate the result in "[]" format and make sure it contains "the score is [Correct]" or "the score is [Incorrect]", for example: “‘ The answer is correct and equivalent to the standard answer, the score is [Correct] ”’ or “‘ The answer is incorrect and does not match the standard answer, the score is [Incorrect] ”’ ## Question: (instruction) ## Standard Answer: (targets) ## Model’s Answer: (answer_to_be_judged)
+Table 6: System Prompts for Multimodal Interaction.

### v3

@@
 Abstract
-Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens.
-Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis.
+Recent studies on end-to-end (E2E) speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens.
+Existing E2E approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis.
 (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation.
 This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations.
-Whereas current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz.
-Experimental results on Spoken Question Answering benchmarks demonstrate that DrVoice establishes new state-of-the-art (SOTA) performance among similar size speech foundation models with relative small amount of data.
+Notably, while current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz, significantly reducing computational cost and alleviating the frequency discrepancy between speech and text tokens and in turn better exploiting LLMs’ capabilities.
+Experimental results demonstrate that DrVoice-7B establishes new state-of-the-art (SOTA) on OpenAudioBench and Big Bench Audio benchmarks, while achieving performance comparable to the SOTA on VoiceBench and UltraEval-Audio benchmarks, making it a leading open-source speech foundation model in $\sim$7B models.
 1 Introduction
-Recent advancements in Large Language Models (LLMs) (OpenAI, 2023; Yang et al., 2024) have demonstrated remarkable performance across various NLP tasks, sparking growing interest in extending their capabilities to multimodal domains encompassing vision and audio.
 Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges.
-The release of GPT-4o (OpenAI, 2024b) is a significant milestone in this field, showing the potential of LLM-based spoken dialogue systems for seamless and natural interactions with users.
+Recently, Large Language Model (LLM) based spoken dialogue systems, exemplified by systems like GPT-4o (OpenAI, 2024b), demonstrate great potential for seamless and natural interactions with users.
 LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs.
 Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules.
-Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., tone, emotion), and significant latency.
+Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., emotion), and significant latency.
 Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024).
@@
 E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
-Recent research efforts on E2E models to address the above challenges have focused on two primary directions Chen et al.
-(2024): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models.
+Recent works on E2E models to address these challenges have focused on two primary directions (Chen et al., 2024a): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models.
 Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation.
 In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously.
-The key distinction lies in whether the generation of speech tokens can influence the subsequent generation of text tokens within the LLM.
-This feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models.
-Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024; KimiTeam et al., 2025).
+The key distinction lies in whether speech token generation can influence the subsequent generation of text tokens within the LLM: this feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models.
+Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024a; KimiTeam et al., 2025).
 Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM.
-While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations.
+While both Text-Driven Speech Models and Joint Speech-Text Models explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations.
 The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins.
-This prevents the LLM from conditioning its textual output on the generated acoustic tokens, thus limiting its ability to control fine-grained paralinguistic attributes like emotion and prosody.
-Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge.
-Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials Chen et al.
-(2024); hence, in this work, we focus on enhancing joint speech-text modeling.
-Recently, Kimi-Audio KimiTeam et al.
-(2025) has set a new state-of-the-art (SOTA) in joint speech-text modeling.
-However, this approach faces notable challenges.
-It not only requires extensive data for training but also incurs significant computational costs due to its 12.5Hz audio representation.
-Furthermore, this high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
-In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR), which compresses the input down to 5Hz.
-DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens.
-For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, which better match the pace of text.
+This prevents the LLM from conditioning its textual output on the generated speech tokens, thus limiting its ability to explore fine-grained paralinguistic attributes like emotion and prosody.
+On the other hand, Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge.
+Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials (Chen et al., 2024a); hence, in this work, we focus on enhancing joint speech-text modeling.
+Recently, Kimi-Audio (KimiTeam et al., 2025) sets a new state-of-the-art (SOTA) in joint speech-text modeling.
+However, this approach still suffers from notable limitations.
+It not only requires extensive training data but also incurs significant computational costs due to its 12.5Hz audio representation.
+Furthermore, the high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024a), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
+In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR).
+For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, effectively alleviating the temporal resolution discrepancy between speech and text tokens.
 During generation, combined speech-text embeddings form the assistant’s autoregressive input.
-The hidden states captured from shared LLM layers are then passed in parallel to a text head for text token prediction, and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens.
-Due to input-output resolution discrepancies, each LLM inference triggers $k$ parallel SRH executions, where $k$ is the grouping factor.
-To further enhance the model’s reasoning and the coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy.
-This approach prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation.
-This intermediate reasoning step serves to improve modality alignment and reduce errors.
-For practical applications, we design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output.
-Experimental results demonstrate the efficacy of CoM in significantly improving performance.
-For reproducibility, we will release all source code and model checkpoints.
+The hidden states captured from shared LLM layer are then passed in parallel to a Text Head for text token prediction and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens.
+To further enhance the model’s reasoning and coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy.
+CoM prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation.
+This intermediate reasoning step improves modality alignment and reduces errors.
+We design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output.
+We also develop a Core-Cocktail training strategy to refine model optimization and LLM knowledge retention.
 Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR).
 This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities.
-2) We introduce a CoM-Mixing training strategy which acts as a curriculum, using the structured reasoning of CoM to scaffold directly speech generation task.
-3) Through comprehensive experiments, we demonstrate that DrVoice achieves new state-of-the-art (SOTA) performance on multiple spoken question answering benchmarks, validating the effectiveness of our proposed architecture and strategies.
+2) We introduce two new training strategy, including a CoM-Mixing training strategy acting as a curriculum, using the structured CoM reasoning to scaffold speech generation, and a Core-Cocktail training strategy for retaining the knowledge of LLMs.
+3) Our comprehensive experimental results reveal that DrVoice-7B achieves new SOTA performance on prominent benchmarks like OpenAudioBench (audio understanding) and Big Bench Audio (reasoning and understanding capabilities of audio-processing models), alongside performance on par with the SOTA on other benchmarks such as VoiceBench (benchmarking LLM-based voice assistants) and UltraEval-Audio (both speech understanding and generation), solidifying its position as a premier open-source speech foundation model among $\sim$7B models.
 2 Related Work
@@
 Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation.
-In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021).
+In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as S3Tokenizer using in CosyVoice (Du et al., 2024a, b, 2025), and prioritized for linguistic content (Hsu et al., 2021).
 While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023).
-In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to their competitive semantic capabilities and compatibility with a more lightweight, diffusion-free vocoder for direct speech synthesis.
+In this work, we use S3Tokenizer as the speech tokenizer due to its solid semantic capabilities and compatibility with the strong Speech Detokenizer from CosyVoice for synthesis.
+S3Tokenizer is a supervised semantic speech tokenizer based on the pre-trained SenseVoice-Large model (An et al., 2024) that enhances the semantic relationship of extracted tokens.
+S3Tokenizer is robust to data noise, and reduces the reliance on clean data.
 E2E Speech Foundation Models.
-Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models.
-Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech.
-Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production.
+Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adapter, LLM, and a streaming speech decoder, and can simultaneously generate text and speech.
+Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker handling speech token production.
 Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
 Joint Speech-Text Models explore two different architectures, Interleaved and Parallel.
-Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024; KimiTeam et al., 2025) conduct parallel decoding.
-The interleaved model VITA-Audio Long et al.
-(2025) achieves efficient audio generation by introducing multiple cross-modal token prediction (MCTP) modules.
-It demonstrates superior performance across multiple benchmarks while maintaining high-quality speech synthesis.
+Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025; Wu et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024a; KimiTeam et al., 2025) conduct parallel decoding.
 The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations.
-To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper Radford et al.
-(2023) features, preserving both semantic and acoustic information.
+To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduces dual-tokenizer combining discrete semantic tokens with continuous Whisper (Radford et al., 2023) features, preserving both semantic and acoustic information.
 In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
@@
 Figure 1 illustrates the architecture of DrVoice.
-The system consists of three main components: (1) Speech Tokenizer and Speech Encoder are designed to process the wave of speech into hidden representation, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, text head, and speech refined head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens.
+The system consists of three main components: (1) Speech Encoder and Speech Tokenizer process the speech wave into hidden representations for the user end and the assistant end respectively, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, a Text Head, and a Speech Refined Head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens.
 The system operates through multimodal input encoding and coordinated speech-text output generation.
 During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head.
-To effectively train this system, we introduce the CoM-Mixing Training strategy.
+To effectively train this system, we introduce the CoM-Mixing Training and Core-Cocktail Training strategies.
 3.1 Speech Tokenization and Detokenization
-Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length.
-For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency.
-Inspired by Kimi-Audio KimiTeam et al.
-(2025) and aiming for improved audio understanding, we utilize the Whisper-Large-v3 Speech Encoder to extract continuous audio representations.
+Aiming for enhanced audio understanding, we utilize Whisper-Large-v3 Speech Encoder to extract continuous audio representations at the user end.
 Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM.
-Finally, the processed continuous representations are stacked with discrete ones and fed into the LLM.
+Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ S3Tokenizer (Du et al., 2024a, b, 2025) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$ at the assistant end, where $T$ is the speech token sequence length.
+For speech detokenization, conditioned on speaker embeddings capturing acoustic details such as timbre, the Flow Matching model (Lipman et al., 2023) converts speech tokens $\mathbf{S}$ into the Mel spectrum for a given speaker.
+Finally, a pre-trained vocoder HiFi-GAN (Kong et al., 2020) transforms the Mel spectrum back into audio signal.
 Figure 1: Overview of DrVoice.
 User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction.
+The MLLM consists of Shared LLM Layer, a Text Head, and a Speech Refined Head (SRH) for token generation.
 The generated speech tokens are then converted to speech waveform by the speech detokenizer.
-Note that SRH through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.
+Note that SRH generates $k$ speech tokens through $k$ autoregressive forward passes, where $k$ is the grouping factor.
+3.2 Multimodal Large Language Model (MLLM)
+Built upon text-LLMs, the MLLM learns joint speech-text modeling to process speech or text inputs while producing parallel speech and text outputs.
+Parallel Joint Speech-Text Model.
+Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold.
+We focus on performing modality alignment exclusively at the assistant end.
+This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
+Leveraging the autoregressive generation capability of LLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the shared LLM layer at each timestep.
+Their embeddings are added as model input, forming a parallel speech-text architecture.
+Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
+$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
+where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively.
+Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with a special token <|SIL|>to align them for each utterance.
+The autoregressive generation process is as follows:
+$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
+where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$.
+This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework.
+To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological designs are introduced for intermodal coordination, including Speech Token Grouping and Ungrouping, and Speech Refined Head.
+Speech Token Grouping.
+A crucial parameter for discrete tokens is the sampling rate, which determines the input/output sequence length.
+GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz.
+Hence, our work adopts 25Hz sampling rate.
+To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024a), a grouping mechanism is designed:
+$\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
+where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates.
+This design compresses the speech token length from $T$ to $T/k$ and the resulting grouped speech representations align better with text.
+Notably, different from Chen et al.
+(2024a) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
+Speech Refined Head (SRH).
+SRH is proposed to enhance speech generation capabilities.
+It utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens.
+While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024a; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details.
+To address this limitation, DrVoice performs an ungrouping process as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
+$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[SLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (4)
+and time splitting
+$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}],$ (5)
+where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$.
+The resulting $\mathbf{H}$ serves as the conditional input for SRH that autoregressively generates speech tokens.
+Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated.
+Speech token prediction is trained to maximize the conditional probability:
+$\mathcal{L}_{\text{SRH}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}_{<i}),$ (6)
+where $s_{i}$ represents the $i$-th speech token.
+By optimizing this objective, SRH learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from SLLM.
+This design enables SRH to effectively leverage the semantic and acoustic information encoded in the hidden representations of SLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
+The E2E training objective $\mathcal{L}_{\text{MLLM}}$ integrates the two losses through multi-task learning:
+$\mathcal{L}_{\text{MLLM}}=\lambda\mathcal{L}_{\text{TH}}+\mu\mathcal{L}_{\text{SRH}},$ (7)
+$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
+where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over the text head, $\lambda$ and $\mu$ are hyperparameters.
+3.3 Training Strategy
+Initialization.
+The Speech Encoder is initialized with the weights of Whisper-Large-v3, while the Shared LLM Layer is initialized using Qwen2.5.
+Additionally, the pre-trained Speech Tokenizer and Detokenizer from CosyVoice are employed and kept frozen throughout the entire training process.
+We initialize SRH with a pre-trained TTS model.
+CoM-Mixing Training.
+The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
+Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs.
+Seven interaction patterns for diverse input-output requirements are summarized in Table 1.
+System prompts are employed to regulate the model’s output behavior.
+For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs.
+Detailed prompts are shown in Appendix B.
+During inference, specifying these system prompts enables the generation of desired output results.
+For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model.
+During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement.
+This flexibility ensures adaptability to varying modality generation demands.
 Table 1: Multimodal Interaction Patterns.
@@
 Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
-3.2 Multimodal Large Language Model (MLLM)
-Built upon text-LLMs, the MLLM learns joint speech-text modeling to process speech or text inputs while producing parallel speech and text outputs.
-Parallel Joint Speech-Text Model.
-Leveraging the autoregressive generation capability of LLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the shared LLM layer at each timestep.
-Their embeddings are added as model input, forming a parallel speech-text architecture.
-Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
-$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
-where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively.
-Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with special token <|SIL|>to align them for each utterance.
-The autoregressive generation process is as follows:
-$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
-where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$.
-This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework.
-To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological components are introduced for intermodal coordination.
-Inner Text Streaming.
-Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold.
-We focus on performing modality alignment exclusively at the assistant end.
-This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
-Speech Token Grouping.
-A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length.
-GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz.
-Hence, our work adopts 25Hz sampling rate.
-To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
-$\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
-where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates.
-This design compresses the speech token length from $T$ to $T/k$ and the grouped speech representations align better with text.
-Different from Chen et al.
-(2024) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
-Speech Refined Head.
-SRH is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens.
-While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details.
-To address this limitation, an ungrouping process is performed as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
-$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[SLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (4)
-and time splitting
-$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}],$ (5)
-where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$, the resulting $\mathbf{H}$ serves as conditional input for SRH that autoregressively generates speech tokens.
-Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated.
-Speech token prediction is trained to maximize the conditional probability:
-$\mathcal{L}_{\text{SRH}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}_{<i}),$ (6)
-where $s_{i}$ represents the $i$-th speech token.
-By optimizing this objective, SRH learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from SLLM.
-This design enables SRH to effectively leverage the semantic and acoustic information encoded in the hidden representations of SLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
-The end-to-end training objective integrates the two losses through multi-task learning paradigm:
-$\mathcal{L}_{\text{MLLM}}=\lambda\mathcal{L}_{\text{TH}}+\mu\mathcal{L}_{\text{SRH}},$ (7)
-$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
-where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over text head, $\lambda$ and $\mu$ are hyperparameters.
-3.3 CoM-Mixing Training strategy
-The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
-Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs.
-Seven interaction patterns for diverse input-output requirements are summarized in Table 1.
-System prompts are employed to regulate the model’s output behavior.
-For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs.
-Detailed prompts are shown in the Appendix A.
-During inference, specifying these system prompts enables the generation of desired output results.
-For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model.
-During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement.
-This flexibility ensures adaptability to varying modality generation demands.
+Core-Cocktail Training.
+We identify a learning rate dilemma: employing a high learning rate significantly compromises the performance of the MLLM, whereas using a low learning rate leads to training stagnation, with the loss decreasing quite slowly.
+To overcome this optimization challenge, we develop a specialized two-stage training strategy, termed “Core-Cocktail”.
+The first stage involves full fine-tuning the entire MLLM with a relatively high learning rate.
+The goal of this phase is to rapidly move the model’s parameters into a more favorable region of the loss landscape.
+However, to counteract the potential performance degradation of the MLLM caused by this aggressive initial training, an intermediate merging step is introduced.
+Drawing inspiration from Xiao et al.
+(2024), the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$).
+This merging creates a new, interpolated model, $M_{r}$, as defined by the following equation:
+$M_{r}\leftarrow\alpha M_{1}+(1-\alpha)M_{0}$
+where $M_{r}$ denotes the merged model and $\alpha$ is the interpolation weight.
+The merging step effectively re-integrates the robust knowledge of the base LLM.
+A smaller $\alpha$ corresponds to a greater preservation of the base LLM’s capabilities.
+The second stage conducts full fine-tuning on the merged model $M_{r}$ using a small learning rate.
+The core-cocktail training approach allows for careful and precise optimization, refining the model’s performance without instability from a high learning rate.
 4 Experiments
-Table 2: Speech-to-Text (S $\rightarrow$ T) and Speech-to-Speech (S $\rightarrow$ S) generation performance on spokenQA benchmarks.
-The results for the S2T baselines are taken from the previous SOTA model Kimi-Audio’s paper (Xu et al., 2025), and the S2S results are taken from the previous SOTA model VITA-Audio’s paper (Long et al., 2025).
-Otherwise, results are taken from their original papers.
-The model size refers only to the parameter count of the LLM backbone, excluding the audio encoder, audio tokenzier, and audio detokenzier module.
-Frame Rate (In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone.
-The average number of text tokens corresponding to one second of speech is denoted by $\tau$.
-Metrics are benchmark-specific (higher is better).
-The Avg.
-score averages two benchmarks.
-The best results in each group are highlighted in bold, while the second-best results are underlined.
-Model Size Frame Rate (In/Out) Llama Q. Web Q. Avg.
-$\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$
+Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best result in each row is in bold).
+With the exception of GLM4-Voice, whose results are cited from Xu et al.
+(2025) and Chen et al.
+(2024b), all other results were generated by running inference on the released checkpoints.
+FR(In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone.
+$\tau$ denotes the average number of text tokens corresponding to one second of speech.
+GLM4-Voice
+MiniCPM
+-o 2.6
+Baichuan
+-Omni-1.5
+Qwen2.5
+-Omni
+Kimi
+-Audio
+Step-Audio2
+-Mini
+DrVoice
 […]
-DrVoice-Small (Our) 2B 5/5 73.0 56.0 42.1 31.1 57.5 43.5
-Moshi (2024) 7B 12.5/12.5 62.3 21.0 26.6 9.2 44.5 15.1
+OpenAudioBench (S2T)
+AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 78.34
+Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 80.33
+Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 57.92
+TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 61.50
+Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 68.10
+Overall 57.70 62.58 64.54 66.34 69.08 60.69 69.24
+VoiceBench (S2T)
+AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.52
+CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 3.77
+SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 68.54
+MMSU 39.75 54.78 57.25 56.38 62.17 52.18 60.31
+OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 79.56
+IFEval 52.80 49.25 54.54 53.88 61.10 38.01 59.30
+AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 98.65
+Overall 59.83 71.69 71.14 72.83 76.93 63.84 76.02
+UltraEval-Audio (S2S)
+AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 49.65
+Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 68.00
+TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 35.35
+Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 37.65
+Overall 42.35 48.05 48.67 50.46 42.79 46.89 47.66
+Big Bench Audio (S2T & S2S)
+S2T 44.8 56.2 47.1 54.2 59.4 50.9 71.6
+S2S 42.7 55.4 44.6 53.6 51.0 47.5 60.9
+Overall 43.8 55.8 45.8 53.9 55.2 49.2 66.3
+Table 3: Speech quality performance on the collections of UltraEval-Audio, in terms of UTMOS for assessing overall speech quality and ASR-WER for assessing alignment between generated speech and text.
+FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech.
 […]
@@
 […]
+Kimi-Audio (2025) 12.5/12.5 3.06 21.06
 […]
-LUCY (S2) (2025) 7B 12.5/48 57.3 51.0 29.3 18.2 43.3 34.6
-[…]
-Kimi-Audio (2025) 8B 12.5/12.5 79.3 - 70.0 - 74.7 -
-[…]
-[…]
-Stream-Omni (2025) 7B 25/25 76.3 65.0 44.2 27.5 60.3 46.3
-[…]
-DrVoice (Our) 8B 5/5 85.0 71.0 65.2 40.1 75.1 55.5
-Table 3: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following Zeng et al.
-(2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text.
-Baseline scores are copied from Zeng et al.
-(2025).
-FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech.
-The best results in each group are bolded.
-[…]
-Mini-Omni (2024a) 0.5B 50/48 3.17 25.28
-DrVoice-Small 2B 5/5 3.87 8.86
-[…]
-Moshi (2025) 7B 12.5/12.5 3.90 7.95
-[…]
-DrVoice 8B 5/5 3.86 6.75
-Table 4: Impact of the Grouping Factor on Llama Q.
-S2T and S2M performance.
-Experiments were conducted on DrVoice-Small w/o.
-CSE, using data subsets for faster exploration.
-The S2T model was trained exclusively on mixed S2T and T2T data, while the S2M model was trained on mixed S2M and T2T data.
-Grouping Factor S2T S2M (T/S)
-1 55.67 4.00 / 2.67
-3 64.67 15.67 / 5.00
-5 63.33 37.67 / 28.00
-7 62.67 36.00 / 16.67
+DrVoice 5/5 4.29 11.2
 Table 4: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
@@
    w/o. CoM-Mixing 58.00 / 49.00 58.00 69.33 / 55.00 68.33 –/– –/– –/–
-Figure 2: Computational Resources across different Grouping Factor.
-Figure 3: Performance Scaling of DrVoice-Small (w/o. Continuous Speech Encoder) on LLaMA Question Benchmark.
 4.1 Experimental Setup
 Datasets.
-Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for pre-training speech refined head.
-17.8K hours of English speech conversations data synthesized from text are adopted for DrVoice training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^1] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b).
-Following the common practice in prior works (Long et al., 2025; KimiTeam et al., 2025; Yao et al., 2024; Zhang et al., 2025), the Speech-to-Speech ($S\rightarrow S$) and Speech-To-Text ($S\rightarrow T$) performance is evaluated on widely used Spoken Question Answering benchmarks, including Web Questions (Berant et al., 2013) and Llama Questions (Nachmani et al., 2024).
+Approximately 100K hours of audio-text paired data for speech-text modality alignment is adopted for pre-training Speech Refined Head.
+For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (Du et al., 2024a, b, 2025), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech.
+Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (Ardila et al., 2020), MELD (Poria et al., 2019), LibriSpeech (Panayotov et al., 2015), SPGISpeech (O’Neill et al., 2021), and Voxpopuli (Wang et al., 2021).
+Following prior works (Yao et al., 2024; KimiTeam et al., 2025; OpenAI, 2024b), we evaluate the performance on the widely used benchmarks, VoiceBench (Chen et al., 2024b) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
 Evaluation Metrics.
-Accuracy is emplyed to assess model performance on datasets.
-The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al.
+Evaluations adhere to the established protocols for each respective benchmark.
+Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (Liu et al., 2023) is used for scoring.
+For AdvBench, the Refusal Rate is reported, while performance on all other benchmarks is assessed with Accuracy.
+The generated speech is transcribed using Whisper-v3-large model (Radford et al., 2022), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text.
+UTMOS (Saeki et al., 2022) is used to evaluate the overall speech quality, following Zeng et al.
 (2025).
 Baselines.
-Representive baselines are selected to cover diverse modeling paradigms: the cascaded system Audio Flamingo 3 (Zhang et al., 2025), Text-Driven models including Freeze-Omni (Défossez et al., 2024), MinMo (Chen et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni2 (Fang et al., 2025), and Qwen2.5-Omni (Xu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025), GLM-4-Voice (Zeng et al., 2025), Stream-Omni (Zhang et al., 2025), and VITA-Audio (Long et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Kimi-Audio (Xu et al., 2025).
-Noting that Kimi-Audio achieves state-of-the-art (SOTA) performance on Speech-to-Text Question Answering, while VITA-Audio sets the SOTA on Speech-to-Speech Question Answering.
-This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
-Implementation Details.
-Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT).
-DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM.
-Grouping factor is set to $k=5$.
-The warmup rate is set to $1\%$ of the total training steps.
-For DrVoice, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for DrVoice-Small and SRH-PT it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$.
-The AdamW method (Loshchilov and Hutter, 2019) is used for optimization.
-Experiments are run on 32x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding.
-For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory.
-It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post training.
-4.2 Results and Discussions
+Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (Yao et al., 2024) and Qwen2.5-Omni (7B) (Xu et al., 2025).
+Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (Zeng et al., 2025), Baichuan-Omni-1.5 (7B) (Li et al., 2025), and Step-Audio2-Mini (8B) (Wu et al., 2025), alongside the parallel model Kimi-Audio (7B) (KimiTeam et al., 2025).
+This suite enables systematic comparisons across mainstream speech-text modeling strategies.
+Implementation Details can be found in Appendix A.
+4.2 Main Results
 […]
-The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (ASR-WER(%)), as shown in Table 3.
-Among models with comparable sizes, DrVoice achieves low ASR-WER, indicating strong alignment when generating both speech and text simultaneously.
-In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a).
-The gap in UTMOS between DrVoice and GLM-4-Voice suggests room for improvement in speech naturalness and intelligibility.
-We conduct extensive ablation study to analyze effects of critical algorithmic designs.
+Computational Efficiency and Speech Quality.
+A key innovation of DrVoice is its remarkable computational efficiency, highlighted in Table 2 and Table 3.
+As shown in the FR (In/Out) rows, DrVoice operates at a frame rate of 5/5, indicating that the LLM backbone processes only 5 audio tokens per second for both input and output.
+This is a substantial reduction compared to other models, which operate at frame rates of 12.5 or 25, thereby significantly lowering computational requirements and potential latency.
+Crucially, this efficiency does not come at the cost of speech quality.
+Table 3 shows that despite its low frame rate, DrVoice produces high-quality and well-aligned speech.
+It achieves UTMOS score of 4.29 for overall speech quality, which is competitive with top-performing models like Qwen2.5-Omni (4.28) and superior to others like Kimi-Audio (3.06).
+With an ASR-WER of 11.2, the model shows robust speech-text alignment and surpasses several baselines.
+Still, it underperforms compared to Qwen2.5-Omni.
+A potential explanation lies in the architectural design: Qwen2.5-Omni’s feeding text directly to its “Talker” module, whereas the proposed model only sends hidden states to SRH.
+Therefore, introducing text to the speech refined head is a logical next step for improving textual control and lowering the ASR-WER.
+This unique combination of SOTA performance, high-quality speech generation, and unparalleled efficiency makes DrVoice a highly powerful and practical model for real-world applications.
+4.3 Ablation and Analyses
+We conduct extensive ablation study and analyses.
+For computational efficiency, some experiments are performed on DrVoice-Small (1.5B).
+More analyses about the data quality and data scaling can be found in Appendix C.
+Core-Cocktail Training Strategy.
+While Stage 1 training causes a performance drop from text baseline of 81.77 to 70.19, Stage 2 reverses this trend, recovering the performance to 74.73.
+This result confirms that the strategy effectively counteracts the initial degradation and leads to an optimized model.
+Further details are available in Appendix C.
 Continuous Speech Encoder.
@@
 In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00.
-This confirms the CSE’s effectiveness in processing and representing speech information.
-Dual-Resolution.
+This confirms the CSE’s effectiveness in representing speech.
+Dual-Resolution Speech Representations (DRSR).
 Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation.
-Grouping Factor.
-Table 4 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency.
+1) Grouping Factor.
 Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M).
@@
 The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5.
-Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, suggesting the efficiency of grouping machinism.
-Speech Refinement.
+Furthermore, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, proving the efficiency of grouping machinism.
+Detail results can be found in Appendix C.
+2) Speech Refinement Head.
 As shown in Table 4, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks.
@@
 Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
+SRH-Pretraining.
 To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model.
@@
 SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%.
-The effect on S2T is smaller (10.2% drop), and it is negligible for T2T.
+The effect on S2T is smaller (10.2% drop), and negligible for T2T.
 This underscores the critical importance of SRH pretraining for refining speech token generation.
@@
 CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
-Data Scaling
-To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o.
-Continous Speech Encoder).
-Each S2S instance can be augmented into 7 distinct multimodal interaction patterns.
-As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements.
-The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
-5 Conclusions and Future Work
+5 Conclusions
 We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations.
-Experimental results demonstrate that DrVoice sets a new state-of-the-art (SOTA) benchmark in both speech-to-speech and speech-to-text generation tasks against current baselines.
+Our experimental results demonstrated that DrVoice establishes a new state-of-the-art (SOTA) on OpenAudioBench for audio understanding and on Big Bench Audio for reasoning capabilities.
+Furthermore, it achieved performance comparable to the SOTA on VoiceBench for voice assistant tasks and on UltraEval-Audio for both speech understanding and generation, confirming its strong and versatile capabilities.
 Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance.
-Our future work will address three key areas: (1) Enhancing Speech Generation Quality: To overcome the limitations of the current suboptimal tokenization and further improve speech fidelity, we plan to explore advanced speech tokenizers, such as the one utilized in CosyVoice.
-This will enable the model to capture finer acoustic details and generate more natural-sounding audio.
-(2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations.
-Inspired by recent advancements like Parrot (Wang et al., 2025), we will investigate the use of a time-division multiplexing (TDM) input stream.
-This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue.
-(3) Expanding to General Audio and Multimodality: Finally, we aim to extend our research beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds.
-The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.
-Appendix A Prompt Templates
+We discuss limitations and future work in Appendix D.
+Appendix A Implementation Details
+Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT).
+The maximum sequence length is set to 2K tokens, which corresponds to an audio duration of approximately 6.8 minutes.
+DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM.
+Grouping factor is set to $k=5$.
+Core-cocktail interpolated factor is set to $\alpha=0$ for extremely maintaining the base LLM capability.
+Multiple loss hyperparameters are set to $\lambda=1$ and $\mu=1$.
+The warmup rate is set to $2\%$ of the total training steps.
+For the two-stage training of DrVoice, the learning rate is decayed from $1\times 10^{-4}$ to $1\times 10^{-5}$ in stage one, and subsequently from $2\times 10^{-5}$ to $2\times 10^{-6}$ in stage two, with both stages utilizing a cosine annealing schedule.
+In contrast, DrVoice-Small and SRH-PT are trained in a single stage, which adopts the same learning rate schedule as the first stage of DrVoice.
+The AdamW method (Loshchilov and Hutter, 2019) is used for optimization.
+Experiments are run on 64x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding.
+For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory.
+It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post-training.
+Appendix B Prompt Templates
+System prompts for multimodal interaction patterns.
 As shown in Table 5, there are five types of system prompts categorized based on the model’s output partitioning.
 During inference, specifying these system prompts enables the generation of desired output results.
+Table 5: System Prompts for Multimodal Interaction.
 Pattern Abbr. System Prompts
@@
 SUC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, and then think of both appropriate text and speech responses at the same time.
-Table 6: System Prompts for Multimodal Interaction.
+Table 6: Training Strategy Performance comparison on the Voicebench benchmark with different training datasets.
+Description AlpacaEval CommonEval SD-QA MMSU OpenBookQA IFEval AdvBench Avg
+Qwen2.5-7B-Instruct (T2T) 4.67 4.34 76.13 69.97 82.20 65.45 99.04 81.86
+Inhouse v1
+OnlyText-Training (T2T) 4.15 3.57 60.15 54.50 66.50 50.22 99.40 69.31
+Stage 2 (S2T) 4.08 3.54 55.88 52.80 69.45 38.48 99.23 66.89
+Inhouse v2
+OnlyText-Training (T2T) 4.22 3.60 68.35 58.10 68.13 65.69 99.62 73.76
+Stage 2 (S2T) 4.10 3.15 64.01 58.30 79.56 55.25 98.27 71.48
+Inhouse v2, MagpiePro, InfGen
+OnlyText-Training (T2T) 4.64 4.26 71.79 69.97 84.40 68.84 99.42 81.77
+Stage 1 (S2T) 4.25 3.09 57.32 54.29 75.82 58.07 99.04 70.19
+Stage 2 (S2T) 4.54 3.35 64.56 61.61 80.44 59.83 98.85 74.73
+Stage 2 w/. ASR data (S2T) 4.52 3.77 68.54 60.31 79.56 59.30 98.65 76.02
+Appendix C More Ablation and Analyses
+Core-Cocktail Training Strategy.
+Table 6 shows the precise trade-offs and benefits of our two-stage approach.
+For instance, under the comprehensive Inhouse v2, MagpiePro, InfGen dataset setting, the model’s average performance after Stage 1 drops to 70.19, a significant decline of over 11 points compared to the text-only baseline (81.77).
+This observation perfectly aligns with our initial hypothesis: the aggressive full fine-tuning with a high learning rate in Stage 1, while intended to rapidly move parameters into a more favorable region of the loss landscape, temporarily compromises the model’s general capabilities, as predicted.
+This performance degradation is precisely the problem that the subsequent steps are designed to rectify.
+Following the intermediate merging step—which “cocktails” the aggressively trained model with the base LLM to re-integrate its robust knowledge—Stage 2 proceeds with careful fine-tuning using a small learning rate.
+As shown in the table, this second stage successfully recovers and refines the model, boosting the average score significantly to 74.73.
+This substantial improvement from Stage 1 demonstrates that our strategy effectively mitigates the instability caused by the initial high-learning-rate training.
+This two-stage process confirms that the Core-Cocktail strategy provides a powerful solution to the learning rate dilemma, enabling rapid adaptation without sacrificing final performance.
+Figure 2: Computational Resources under 17K hours training data across different Grouping Factor.
+Data Quality.
+The OnlyText-Training (T2T) results in Table 6 serve as a powerful indicator of the maximum potential embedded within the textual content of our datasets.
+There is a clear correlation between data quality and the final model’s performance: as the T2T average score improves from Inhouse v1 (69.31) to our most comprehensive dataset (81.77), the final Stage 2 (S2T) model performance also sees a substantial lift.
+Notably, our final dataset mixture enables the T2T model to reach an average score of 81.77, almost perfectly matching the text-only Qwen2.5-7B-Instruct backbone.
+This confirms the exceptional quality of our instructional data and establishes a high-performance ceiling.
+Our final S2T model (74.73) successfully translates a significant portion of this potential into the multimodal domain.
+Furthermore, the results underscore the critical importance of data composition for specific tasks.
+The CommonEval benchmark, which evaluates models on real human speech, is a case in point.
+Our standard Stage 2 (S2T) model achieves a score of 3.35 on this challenging benchmark.
+However, by augmenting the training data with ASR-derived pairs (Stage 2 w/.
+ASR data), the CommonEval score noticeably increases to 3.77.
+This gain is because ASR data exposes the model to the natural disfluencies, varied intonations, and ambient noise present in real-world speech, enhancing its robustness.
+Data Scaling.
+To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o.
+Continous Speech Encoder).
+Each S2S instance can be augmented into 7 distinct multimodal interaction patterns.
+As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements.
+The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
+Figure 3: Performance Scaling of DrVoice-Small (w/o. Continuous Speech Encoder) on LLaMA Question Benchmark.
+Table 7: Impact of the Grouping Factor on Llama Q.
+S2T and S2M performance, using DrVoice-Small (1.5B) w/o CSE and data subsets for faster experiments.
+The S2T model is trained exclusively on mixed S2T and T2T data, while the S2M model is trained on mixed S2M and T2T data.
+Grouping Factor S2T S2M (T/S)
+1 55.67 4.00 / 2.67
+3 64.67 15.67 / 5.00
+5 63.33 37.67 / 28.00
+7 62.67 36.00 / 16.67
+Grouping Factor.
+Table 7 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency.
+Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M).
+For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain).
+The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5.
+Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, proving the efficiency of grouping machinism.
+Retention vs Forgetting.
+An attempt is made to select suitable open-source datasets for speech synthesis.
+All results are from fine-tuning Qwen2.5-7B-Instruct at a low learning rate (1e-5), so they primarily indicate retention vs.
+forgetting rather than absolute capability gains.
+The gap to the baseline quantifies forgetting: magpie_pro_llama3_3_500k_filtered and InfGen exhibit the smallest overall deltas (about 3–4 points on both suites), while very large, noisier corpora like Inf7M induce substantial forgetting despite their size.
+Quality and filtering matter more than scale: openorca_gpt4 forgets less than openorca_gpt3_5 with fewer tokens, and the magpie_pro family maintains instruction-following and factual QA best per token.
+Forgetting is selective: some submetrics improve (e.g., LlamaQ rises for Inf7M/tulu_3_sft; AdvBench reaches 100 for tulu_3_sft), but these gains often trade off against instruction and QA scores.
+Overall, compact, well-filtered, model-proximal data minimizes destructive drift at 1e-5, whereas bulk token count tends to amplify forgetting, so the most stable recipes preserve the base model while making targeted adjustments.
+Table 8: Benchmark Results: VoiceBench (T2T)
+Data VoiceBench
+AlpacaEval CommonEval SD-QA MMSU OBQA IFEval AdvBench Overall
+Qwen2.5-7B-Instruct 4.67 4.34 76.13 69.97 82.20 65.45 99.04 81.86
+Inf7M 3.83 3.50 69.98 55.43 71.87 48.79 94.42 69.58
+InfGen 4.51 4.13 74.14 62.46 76.04 58.35 98.46 77.46
+openorca_gpt4 4.33 4.03 73.96 61.00 73.85 54.05 96.35 75.20
+openorca_gpt3_5 4.35 4.17 71.07 57.84 66.59 49.39 95.00 72.90
+Evol_Instruct_Code 4.19 4.06 69.26 63.14 81.54 56.40 90.38 75.10
+Evol_Instruct 4.09 3.91 71.61 60.38 72.31 54.02 98.27 73.80
+magpie_pro_llama3_1_300k 4.42 4.08 72.69 58.33 74.95 56.08 86.73 74.11
+magpie_pro_mt_llama3_1_300k 4.46 4.05 73.60 58.07 60.00 60.05 86.35 72.61
+numinamath_cot 4.11 3.78 70.16 55.43 68.35 52.00 97.69 71.63
+OpenHermes_v2_5 4.05 3.71 71.61 62.20 74.07 51.59 82.88 71.08
+Synthia_v1_3 4.30 3.96 72.33 63.63 76.70 53.32 97.69 75.55
+tulu_3_sft 4.04 3.73 65.64 61.74 79.78 61.78 100.00 74.91
+magpie_pro_llama3_3_500k_filtered 4.48 4.17 73.42 64.31 81.32 64.81 92.88 78.53
+Table 9: Benchmark Results: UltraEval-Audio (T2T) & OpenAudioBench (T2T)
+Data Tokens (M) UltraEval-Audio OpenAudioBench Total
+AlpacaEval LlamaQ TriviaQA WebQ Overall Reasoning QA Avg
+Qwen2.5-7B-Instruct 81.41 82.33 53.91 52.56 67.55 58.00 75.10
+Inf7M 2406 55.35 84.67 46.29 47.59 58.48 48.51 64.13
+InfGen 1022.4 74.29 82.33 55.96 51.18 65.94 57.43 71.95
+openorca_gpt4 361.8 66.21 83.00 54.39 49.56 63.29 54.46 69.50
+openorca_gpt3_5 1101 55.91 83.00 54.69 48.47 60.52 48.51 66.74
+Evol_Instruct_Code 28.3 52.32 81.33 52.64 44.24 57.63 52.48 67.39
+Evol_Instruct 69.2 56.62 81.67 53.81 47.15 59.81 53.47 67.44
+magpie_pro_llama3_1_300k 213.9 76.16 83.67 57.91 55.02 68.19 53.96 70.46
+magpie_pro_mt_llama3_1_300k 323.2 75.35 83.00 54.00 54.53 66.72 57.43 69.38
+numinamath_cot 463.2 64.04 80.67 54.88 50.64 62.56 50.00 66.81
+OpenHermes_v2_5 375.9 58.54 83.67 54.49 49.85 61.64 54.46 66.55
+Synthia_v1_3 59 61.67 83.67 53.81 47.88 61.76 55.45 69.28
+tulu_3_sft 570.2 55.91 84.33 53.52 49.41 60.79 55.94 68.62
+magpie_pro_llama3_3_500k_filtered 314.6 78.64 79.00 57.32 54.18 67.29 51.49 72.53
+Appendix D Limitations and Future Work
+Our future work will address limitations of this work and advance three key areas: (1) Enhancing Speech Generation Quality: The model currently exhibits weaker alignment between text and speech in speech-to-speech generation when compared to Qwen2.5-Omni.
+This performance gap can be attributed to a key architectural difference in the final speech generation stage.
+The Qwen2.5-Omni method directly feeds text into its talker module to ensure high textual fidelity.
+In contrast, DrVoice only inputs hidden states into its speech refined head (SRH).
+Therefore, a promising direction for future work is to incorporate text as an additional input to the SRH.
+It is hypothesized that this modification will provide more explicit textual guidance, further enhancing speech-text alignment and leading to a significant reduction in ASR-WER.
+(2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations.
+Inspired by recent advancements like Parrot (Wang et al., 2025), future work will investigate the use of a time-division multiplexing (TDM) input stream.
+This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue.
+(3) Expanding to General Audio and Multimodality: Finally, a promising avenue is to extend the model’s capabilities beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds.
+The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.

### v4

@@
 Notably, while current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz, significantly reducing computational cost and alleviating the frequency discrepancy between speech and text tokens and in turn better exploiting LLMs’ capabilities.
-Experimental results demonstrate that DrVoice-7B establishes new state-of-the-art (SOTA) on OpenAudioBench and Big Bench Audio benchmarks, while achieving performance comparable to the SOTA on VoiceBench and UltraEval-Audio benchmarks, making it a leading open-source speech foundation model in $\sim$7B models.
+Experimental results demonstrate that DrVoice-7B establishes new state-of-the-art (SOTA) on prominent speech benchmarks including OpenAudioBench, VoiceBench, UltraEval-Audio and Big Bench Audio, making it a leading open-source speech foundation model in $\sim$7B models.
 1 Introduction
 Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges.
-Recently, Large Language Model (LLM) based spoken dialogue systems, exemplified by systems like GPT-4o (OpenAI, 2024b), demonstrate great potential for seamless and natural interactions with users.
+Recently, Large Language Model (LLM) based spoken dialogue systems, exemplified by systems like GPT-4o (openai2024hello), demonstrate great potential for seamless and natural interactions with users.
 LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs.
@@
 Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., emotion), and significant latency.
-Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024).
+Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (DBLP:journals/corr/abs-2311-07919; DBLP:journals/corr/abs-2407-10759).
 E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs.
@@
 E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
-Recent works on E2E models to address these challenges have focused on two primary directions (Chen et al., 2024a): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models.
+Recent works on E2E models to address these challenges have focused on two primary directions (chen2024slam): Text-Driven Speech Models (yao2024minicpm; li2025baichuan) and Joint Speech-Text Models.
 Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation.
@@
 The key distinction lies in whether speech token generation can influence the subsequent generation of text tokens within the LLM: this feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models.
-Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024a; KimiTeam et al., 2025).
+Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (DBLP:conf/iclr/ZengDLZJD025; DBLP:journals/corr/abs-2410-17799) and parallel speech-text modeling (DBLP:journals/corr/abs-2410-00037; chen2024slam; kimiteam2025kimiaudiotechnicalreport).
 Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM.
@@
 On the other hand, Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge.
-Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials (Chen et al., 2024a); hence, in this work, we focus on enhancing joint speech-text modeling.
-Recently, Kimi-Audio (KimiTeam et al., 2025) sets a new state-of-the-art (SOTA) in joint speech-text modeling.
+Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials (chen2024slam); hence, in this work, we focus on enhancing joint speech-text modeling.
+Recently, Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) sets a new state-of-the-art (SOTA) in joint speech-text modeling.
 However, this approach still suffers from notable limitations.
 It not only requires extensive training data but also incurs significant computational costs due to its 12.5Hz audio representation.
-Furthermore, the high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024a), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
+Furthermore, the high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (chen2024slam), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
 In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR).
@@
 The hidden states captured from shared LLM layer are then passed in parallel to a Text Head for text token prediction and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens.
-To further enhance the model’s reasoning and coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy.
+To further enhance the model’s reasoning and coherence of its output, we incorporate a Chain-of-Modality (CoM) (DBLP:conf/emnlp/ZhangLZZWZQ23) strategy.
 CoM prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation.
@@
 Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR).
-This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities.
+This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs (empirically achieving a reduction of nearly 50% in GPU hours for training) and better preserving the LLM’s semantic processing capabilities.
 2) We introduce two new training strategy, including a CoM-Mixing training strategy acting as a curriculum, using the structured CoM reasoning to scaffold speech generation, and a Core-Cocktail training strategy for retaining the knowledge of LLMs.
-3) Our comprehensive experimental results reveal that DrVoice-7B achieves new SOTA performance on prominent benchmarks like OpenAudioBench (audio understanding) and Big Bench Audio (reasoning and understanding capabilities of audio-processing models), alongside performance on par with the SOTA on other benchmarks such as VoiceBench (benchmarking LLM-based voice assistants) and UltraEval-Audio (both speech understanding and generation), solidifying its position as a premier open-source speech foundation model among $\sim$7B models.
+3) Our comprehensive experimental results reveal that DrVoice-7B achieves new SOTA performance on prominent benchmarks including OpenAudioBench (audio understanding), VoiceBench (benchmarking LLM-based voice assistants), UltraEval-Audio (both speech understanding and generation) and Big Bench Audio (reasoning and understanding capabilities of audio-processing models), solidifying its position as a premier open-source speech foundation model among $\sim$7B models.
 2 Related Work
 Speech Tokenization.
-Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations.
+Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (radford2022whisper), and discrete representations.
 Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation.
-In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as S3Tokenizer using in CosyVoice (Du et al., 2024a, b, 2025), and prioritized for linguistic content (Hsu et al., 2021).
-While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023).
+In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (DBLP:journals/tmlr/DefossezCSA23), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (DBLP:conf/nips/HassidRNGCKCDSD23) or supervised learning as S3Tokenizer using in CosyVoice (DBLP:journals/corr/abs-2407-05407; DBLP:journals/corr/abs-2412-10117; DBLP:journals/corr/abs-2505-17589), and prioritized for linguistic content (DBLP:journals/taslp/HsuBTLSM21).
+While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (DBLP:journals/corr/abs-2308-16692), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (DBLP:conf/iclr/ZengDLZJD025; DBLP:journals/corr/abs-2410-08035; DBLP:conf/emnlp/ZhangLZZWZQ23; DBLP:journals/taslp/BorsosMVKPSRTGTZ23).
 In this work, we use S3Tokenizer as the speech tokenizer due to its solid semantic capabilities and compatibility with the strong Speech Detokenizer from CosyVoice for synthesis.
-S3Tokenizer is a supervised semantic speech tokenizer based on the pre-trained SenseVoice-Large model (An et al., 2024) that enhances the semantic relationship of extracted tokens.
+S3Tokenizer is a supervised semantic speech tokenizer based on the pre-trained SenseVoice-Large model (DBLP:journals/corr/abs-2407-04051) that enhances the semantic relationship of extracted tokens.
 S3Tokenizer is robust to data noise, and reduces the reliance on clean data.
 E2E Speech Foundation Models.
-Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adapter, LLM, and a streaming speech decoder, and can simultaneously generate text and speech.
-Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker handling speech token production.
+Text-Driven Speech Models (DBLP:journals/corr/abs-2409-06666; DBLP:journals/corr/abs-2411-00774; DBLP:journals/corr/abs-2501-01957; yao2024minicpm; huang2025stepaudiounifiedunderstandinggeneration; DBLP:journals/corr/abs-2501-06282) integrate speech encoder, adapter, LLM, and a streaming speech decoder, and can simultaneously generate text and speech.
+Qwen2.5-Omni (DBLP:journals/corr/abs-2503-20215) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker handling speech token production.
 Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
 Joint Speech-Text Models explore two different architectures, Interleaved and Parallel.
-Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025; Wu et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024a; KimiTeam et al., 2025) conduct parallel decoding.
-The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations.
-To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduces dual-tokenizer combining discrete semantic tokens with continuous Whisper (Radford et al., 2023) features, preserving both semantic and acoustic information.
+Interleaved models (DBLP:journals/corr/abs-2410-17799; DBLP:conf/iclr/ZengDLZJD025; li2025baichuan; DBLP:journals/corr/abs-2505-03739; DBLP:journals/corr/abs-2507-16632) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (DBLP:journals/corr/abs-2410-00037; DBLP:journals/corr/abs-2408-16725; DBLP:journals/corr/abs-2410-11190; chen2024slam; kimiteam2025kimiaudiotechnicalreport) conduct parallel decoding.
+The parallel model Moshi (DBLP:journals/corr/abs-2410-00037) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations.
+To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) introduces dual-tokenizer combining discrete semantic tokens with continuous Whisper (DBLP:conf/icml/RadfordKXBMS23) features, preserving both semantic and acoustic information.
 In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
@@
 Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM.
-Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ S3Tokenizer (Du et al., 2024a, b, 2025) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$ at the assistant end, where $T$ is the speech token sequence length.
-For speech detokenization, conditioned on speaker embeddings capturing acoustic details such as timbre, the Flow Matching model (Lipman et al., 2023) converts speech tokens $\mathbf{S}$ into the Mel spectrum for a given speaker.
-Finally, a pre-trained vocoder HiFi-GAN (Kong et al., 2020) transforms the Mel spectrum back into audio signal.
+Semantic tokens have been widely used for speech tokenization (DBLP:conf/emnlp/ZhangLZZWZQ23; DBLP:journals/taslp/BorsosMVKPSRTGTZ23), due to their strong alignment with text (DBLP:journals/corr/abs-2308-16692); therefore, we employ S3Tokenizer (DBLP:journals/corr/abs-2407-05407; DBLP:journals/corr/abs-2412-10117; DBLP:journals/corr/abs-2505-17589) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$ at the assistant end, where $T$ is the speech token sequence length.
+For speech detokenization, conditioned on speaker embeddings capturing acoustic details such as timbre, the Flow Matching model (DBLP:conf/iclr/LipmanCBNL23) converts speech tokens $\mathbf{S}$ into the Mel spectrum for a given speaker.
+Finally, a pre-trained vocoder HiFi-GAN (DBLP:conf/nips/KongKB20) transforms the Mel spectrum back into audio signal.
 Figure 1: Overview of DrVoice.
@@
 Parallel Joint Speech-Text Model.
-Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold.
+Inspired by Moshi (kyutai2024moshi), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold.
 We focus on performing modality alignment exclusively at the assistant end.
@@
 A crucial parameter for discrete tokens is the sampling rate, which determines the input/output sequence length.
-GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz.
+GLM-4-Voice (DBLP:conf/iclr/ZengDLZJD025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz.
 Hence, our work adopts 25Hz sampling rate.
-To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024a), a grouping mechanism is designed:
+To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (chen2024slam), a grouping mechanism is designed:
 $\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
@@
 This design compresses the speech token length from $T$ to $T/k$ and the resulting grouped speech representations align better with text.
-Notably, different from Chen et al.
-(2024a) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
+Notably, different from chen2024slam that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
 Speech Refined Head (SRH).
@@
 It utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens.
-While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024a; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details.
+While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (chen2024slam; DBLP:journals/corr/abs-2410-08035), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details.
 To address this limitation, DrVoice performs an ungrouping process as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
@@
 CoM-Mixing Training.
-The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
+The chain-of-modality (CoM) methodology (DBLP:conf/emnlp/ZhangLZZWZQ23) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical.
 Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs.
@@
 However, to counteract the potential performance degradation of the MLLM caused by this aggressive initial training, an intermediate merging step is introduced.
-Drawing inspiration from Xiao et al.
-(2024), the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$).
+Drawing inspiration from DBLP:conf/acl/XiaoLZX24, the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$).
 This merging creates a new, interpolated model, $M_{r}$, as defined by the following equation:
@@
 4 Experiments
-Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best result in each row is in bold).
-With the exception of GLM4-Voice, whose results are cited from Xu et al.
-(2025) and Chen et al.
-(2024b), all other results were generated by running inference on the released checkpoints.
+Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best and second-best results in each row are in bold and underlined, respectively).
+With the exception of GLM4-Voice, whose results are cited from DBLP:journals/corr/abs-2503-20215 and DBLP:journals/corr/abs-2410-17196, and the BBA results, which are cited from  coreteam2025mimoaudio except for MiniCPM-o 2.6 that we reproduce, all other results were generated by running inference on the released checkpoints.
 FR(In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone.
 $\tau$ denotes the average number of text tokens corresponding to one second of speech.
-GLM4-Voice
-MiniCPM
--o 2.6
-Baichuan
--Omni-1.5
-Qwen2.5
--Omni
-Kimi
--Audio
-Step-Audio2
--Mini
-DrVoice
+GLM4-Voice MiniCPM -o 2.6 Baichuan -Omni-1.5 Qwen2.5 -Omni Kimi -Audio Step-Audio2 -Mini DrVoice
 […]
 OpenAudioBench (S2T)
-AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 78.34
-Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 80.33
-Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 57.92
-TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 61.50
-Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 68.10
-Overall 57.70 62.58 64.54 66.34 69.08 60.69 69.24
+AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 82.61
+Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 83.00
+Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 59.90
+TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 64.50
+Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 70.20
+Overall 57.70 62.58 64.54 66.34 69.08 60.69 72.04
 VoiceBench (S2T)
-AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.52
-CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 3.77
-SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 68.54
-MMSU 39.75 54.78 57.25 56.38 62.17 52.18 60.31
-OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 79.56
-IFEval 52.80 49.25 54.54 53.88 61.10 38.01 59.30
-AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 98.65
-Overall 59.83 71.69 71.14 72.83 76.93 63.84 76.02
+AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.74
+CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 4.08
+SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 64.30
+MMSU 39.75 54.78 57.25 56.38 62.17 52.18 67.27
+OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 82.20
+IFEval 52.80 49.25 54.54 53.88 61.10 38.01 71.39
+AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 99.62
+Overall 59.83 71.69 71.14 72.83 76.93 63.84 80.17
 UltraEval-Audio (S2S)
-AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 49.65
-Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 68.00
-TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 35.35
-Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 37.65
-Overall 42.35 48.05 48.67 50.46 42.79 46.89 47.66
+AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 59.29
+Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 75.33
+TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 46.09
+Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 45.92
+Overall 42.35 48.05 48.67 50.46 42.79 46.89 56.66
 Big Bench Audio (S2T & S2S)
-S2T 44.8 56.2 47.1 54.2 59.4 50.9 71.6
-S2S 42.7 55.4 44.6 53.6 51.0 47.5 60.9
-Overall 43.8 55.8 45.8 53.9 55.2 49.2 66.3
+S2T 44.8 56.2 47.1 54.2 59.4 50.9 76.9
+S2S 42.7 55.4 44.6 53.6 51.0 47.5 71.2
+Overall 43.8 55.8 45.8 53.9 55.2 49.2 74.0
 Table 3: Speech quality performance on the collections of UltraEval-Audio, in terms of UTMOS for assessing overall speech quality and ASR-WER for assessing alignment between generated speech and text.
@@
 […]
-Kimi-Audio (2025) 12.5/12.5 3.06 21.06
+Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) 12.5/12.5 3.06 21.06
 […]
-DrVoice 5/5 4.29 11.2
-Table 4: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
+DrVoice 5/5 4.29 8.36
+Table 4: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing on Llama Questions. S2M, S2T, etc are defined in Table 1.
 Model S2M (T/S) S2T T2M (T/S) T2T STC (T/S) SAC (T/S) SUC (T/S)
@@
 Approximately 100K hours of audio-text paired data for speech-text modality alignment is adopted for pre-training Speech Refined Head.
-For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (Du et al., 2024a, b, 2025), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech.
-Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (Ardila et al., 2020), MELD (Poria et al., 2019), LibriSpeech (Panayotov et al., 2015), SPGISpeech (O’Neill et al., 2021), and Voxpopuli (Wang et al., 2021).
-Following prior works (Yao et al., 2024; KimiTeam et al., 2025; OpenAI, 2024b), we evaluate the performance on the widely used benchmarks, VoiceBench (Chen et al., 2024b) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
+For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (DBLP:journals/corr/abs-2407-05407; DBLP:journals/corr/abs-2412-10117; DBLP:journals/corr/abs-2505-17589), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech.
+Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (DBLP:conf/lrec/ArdilaBDKMHMSTW20), MELD (DBLP:conf/acl/PoriaHMNCM19), LibriSpeech (DBLP:conf/icassp/PanayotovCPK15), SPGISpeech (DBLP:conf/interspeech/ONeillLMNZKBDFS21), and Voxpopuli (DBLP:conf/acl/WangRLWTHWPD20).
+Following prior works (yao2024minicpm; kimiteam2025kimiaudiotechnicalreport; openai2024hello), we evaluate the performance on the widely used benchmarks, VoiceBench (DBLP:journals/corr/abs-2410-17196) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
 Evaluation Metrics.
 Evaluations adhere to the established protocols for each respective benchmark.
-Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (Liu et al., 2023) is used for scoring.
+Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (DBLP:conf/emnlp/LiuIXWXZ23) is used for scoring.
 For AdvBench, the Refusal Rate is reported, while performance on all other benchmarks is assessed with Accuracy.
-The generated speech is transcribed using Whisper-v3-large model (Radford et al., 2022), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text.
-UTMOS (Saeki et al., 2022) is used to evaluate the overall speech quality, following Zeng et al.
-(2025).
+The generated speech is transcribed using Whisper-v3-large model (radford2022whisper), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text.
+UTMOS (DBLP:conf/interspeech/SaekiXNKTS22) is used to evaluate the overall speech quality, following DBLP:conf/iclr/ZengDLZJD025.
 Baselines.
-Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (Yao et al., 2024) and Qwen2.5-Omni (7B) (Xu et al., 2025).
-Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (Zeng et al., 2025), Baichuan-Omni-1.5 (7B) (Li et al., 2025), and Step-Audio2-Mini (8B) (Wu et al., 2025), alongside the parallel model Kimi-Audio (7B) (KimiTeam et al., 2025).
+Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (yao2024minicpm) and Qwen2.5-Omni (7B) (DBLP:journals/corr/abs-2503-20215).
+Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (DBLP:conf/iclr/ZengDLZJD025), Baichuan-Omni-1.5 (7B) (li2025baichuan), and Step-Audio2-Mini (8B) (DBLP:journals/corr/abs-2507-16632), alongside the parallel model Kimi-Audio (7B) (kimiteam2025kimiaudiotechnicalreport).
 This suite enables systematic comparisons across mainstream speech-text modeling strategies.
@@
 It achieves UTMOS score of 4.29 for overall speech quality, which is competitive with top-performing models like Qwen2.5-Omni (4.28) and superior to others like Kimi-Audio (3.06).
-With an ASR-WER of 11.2, the model shows robust speech-text alignment and surpasses several baselines.
+With an ASR-WER of 8.36, the model shows robust speech-text alignment and surpasses several baselines.
 Still, it underperforms compared to Qwen2.5-Omni.
 A potential explanation lies in the architectural design: Qwen2.5-Omni’s feeding text directly to its “Talker” module, whereas the proposed model only sends hidden states to SRH.
-Therefore, introducing text to the speech refined head is a logical next step for improving textual control and lowering the ASR-WER.
 This unique combination of SOTA performance, high-quality speech generation, and unparalleled efficiency makes DrVoice a highly powerful and practical model for real-world applications.
@@
 As shown in Table 4, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs.
+In this setting, user audio input is processed solely using the discrete speech tokenizer (same as the assistant’s output tokenizer), without the Whisper features.
 Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation.
@@
 As shown in Table 4, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks.
+To handle the resolution mismatch when SRH is removed, the hidden state from the shared LLM is passed through a projection layer ($d_{model}\to k\times d_{speech\_token}$) and then split into k tokens.
+This setup simulates a standard parallel prediction approach without autoregressive refinement.
 By comparing the model with SRH (w/o.
@@
 We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations.
-Our experimental results demonstrated that DrVoice establishes a new state-of-the-art (SOTA) on OpenAudioBench for audio understanding and on Big Bench Audio for reasoning capabilities.
-Furthermore, it achieved performance comparable to the SOTA on VoiceBench for voice assistant tasks and on UltraEval-Audio for both speech understanding and generation, confirming its strong and versatile capabilities.
-Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance.
+Our experimental results demonstrated that DrVoice establishes a new state-of-the-art (SOTA) on OpenAudioBench for audio understanding, VoiceBench for voice assistant tasks, UltraEval-Audio for both speech understanding and generation and on Big Bench Audio for reasoning capabilities, confirming its strong and versatile capabilities.
+Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency (empirically achieving a reduction of nearly 50% in GPU hours for training) without sacrificing performance.
 We discuss limitations and future work in Appendix D.
+[…]
 Appendix A Implementation Details
@@
 In contrast, DrVoice-Small and SRH-PT are trained in a single stage, which adopts the same learning rate schedule as the first stage of DrVoice.
-The AdamW method (Loshchilov and Hutter, 2019) is used for optimization.
-Experiments are run on 64x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding.
-For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory.
+The AdamW method (DBLP:conf/iclr/LoshchilovH19) is used for optimization.
+Experiments are run on 64x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(DBLP:journals/corr/abs-1905-12322) to accelerate training and decoding.
+For DrVoice, DeepSpeed ZeRO-2 (DBLP:conf/sc/RajbhandariRRH20) is implemented to prevent GPU out-of-memory.
 It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post-training.
@@
 This gain is because ASR data exposes the model to the natural disfluencies, varied intonations, and ambient noise present in real-world speech, enhancing its robustness.
+SRH-Pretraining Policy.
+Two pre-training strategies were employed for the Speech Refined Head (SRH).
+The first strategy involves independently training a Text-to-Speech (TTS) task on the Qwen2.5-0.5B model in a T2M format with 25Hz token rate; the resulting trained weights are then loaded onto the SRH.
+In contrast, the second strategy freezes the Speech Encoder, Adaptor, Shared LLM Layer, and the Text Head, while exclusively training the SRH with a streaming TTS objective.
+Experiments demonstrated that the second strategy yields a substantial improvement in Speech-to-Speech (S2S) performance on the UltraEval-Audio benchmark, with the score increasing from 47.66 to 56.66.
+This significant gain is attributed to the closer alignment between the pre-training task in the second strategy and the format of the downstream task.
+Reinforcement learning.
+To further enhance the model’s capabilities in real-world speech comprehension, instruction following, and reasoning, reinforcement learning (RL) methods were employed.
+Specifically, to address the scarcity of authentic speech data, a preference dataset was constructed using open-source ASR datasets.
+A text LLM was used to identify samples suitable for user queries, for which canonical responses were then generated by another text LLM.
+Subsequently, the DrVoice-SFT model performed inference on the original speech queries to establish the preference data.
+For the instruction-following capability, the open-source preference dataset  [^4] https://huggingface.co/datasets/allenai/tulu-3-pref-personas-instruction-following was utilized.
+Training for these first two objectives was performed using the DPO (DBLP:conf/nips/RafailovSMMEF23) algorithm.
+Finally, to improve reasoning ability, the model underwent GSPO (DBLP:journals/corr/abs-2507-18071) training on mathematical problems  [^5] https://github.com/openai/prm800k.
 Data Scaling.
@@
 Appendix D Limitations and Future Work
-Our future work will address limitations of this work and advance three key areas: (1) Enhancing Speech Generation Quality: The model currently exhibits weaker alignment between text and speech in speech-to-speech generation when compared to Qwen2.5-Omni.
-This performance gap can be attributed to a key architectural difference in the final speech generation stage.
-The Qwen2.5-Omni method directly feeds text into its talker module to ensure high textual fidelity.
-In contrast, DrVoice only inputs hidden states into its speech refined head (SRH).
-Therefore, a promising direction for future work is to incorporate text as an additional input to the SRH.
-It is hypothesized that this modification will provide more explicit textual guidance, further enhancing speech-text alignment and leading to a significant reduction in ASR-WER.
-(2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations.
-Inspired by recent advancements like Parrot (Wang et al., 2025), future work will investigate the use of a time-division multiplexing (TDM) input stream.
+Our future work will address limitations of this work and advance two key areas: (1) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations.
+Inspired by recent advancements like Parrot (wang2025parrot), future work will investigate the use of a time-division multiplexing (TDM) input stream.
 This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue.